### v2

# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

Abstract
A residual-type a posteriori error estimation is developed for an interior penalty virtual element method (IPVEM) to solve a Kirchhoff plate bending problem with inhomogeneous boundary value conditions. The computable error estimator is incorporated. We derive the reliability and efficiency of the a posteriori error bound by constructing an enriching operator and establishing some related error estimates that align with interior penalty finite element methods. As an outcome of the error estimator, an adaptive VEM is introduced by means of the mesh refinement strategy with the one-hanging-node rule. Numerical results on several benchmark tests confirm the robustness of the proposed error estimator and show the efficiency of the resulting adaptive VEM.
keywords:
Kirchhoff plate, Interior penalty virtual element method, A posteriori error analysis, Enriching operator, Adaptive methods.
1 Introduction
Virtual element methods (VEMs) generalize standard finite element methods to accommodate general polytopal meshes. Initially proposed and analyzed in [3], with subsequent foundational work in [1, 4], VEMs offer advantages in handling complex geometries and high-regularity solutions compared to traditional finite element methods. For the plate bending problem or the biharmonic equation, Brezzi and Marini introduced an $H^{2}$-conforming VEM in [11], establishing convergence results in the $H^{2}$ norm. Chinosi and Marini further improved the method in [17] by deriving optimal order error estimates in the $L^{2}$ and $H^{1}$ norms through VEM enhancement techniques. Nonconforming VEMs have also been developed, including the $C^{0}$-continuous $H^{2}$-nonconforming VEM presented in [28] and the Morley-type VEM discussed in [30]. Chen and Huang extended the approach to fully $H^{m}$-nonconforming VEMs for $2m$-th order problems in any dimension in [15]. Recently, the interior penalty virtual element method (IPVEM) has been explored in [29] for the biharmonic equation. Compared with the conforming or nonconforming VEMs for the biharmonic problem, the IPVEM has advantages in reducing the number of degrees of freedom (DoFs), which can be regarded as a combination of the virtual element space and discontinuous Galerkin scheme, since the resulting global discrete space is not $C^{0}$-continuous and an interior penalty formulation is adopted to enforce the continuity of the solution. This method was adapted for the fourth-order singular perturbation problem in [27], with some adaptations to the original space. These modifications encompass alterations in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast, Ref. [18] employs the original IPVEM formulation from [29] to address the fourth-order singular perturbation problem (2.1), incorporating techniques from the modified Morley finite element method as described in [25].
On the other hand, due to the large flexibility of the meshes in VEMs, the a posteriori error analysis of VEMs is well worth studying. For the fourth-order problem, a residual-based a posteriori error estimate for the $C^{1}$-continuous VEM was developed in [16], addressing both the reliability and efficiency of the error bound. Recently, the lowest-order nonconforming or Morley-type virtual element has been applied to the biharmonic equation, with a residual-based a posteriori error estimator derived in [13]. It’s worth noting that the a posteriori error analysis in [13] circumvents any trace of second derivatives by using some computable conforming companion operator, which significantly reduces the computational complexity of the error estimators. This reduction is facilitated by the special property of the interpolation operator for the Morley virtual element (see Lemma 5.1).
In this paper, we consider the a posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates. We provide a detailed description for implementing the IPVEM with inhomogeneous boundary conditions. For the sake of the construction of error estimators, unlike the previous interior penalty formulations in [29, 18], we substitute the $H^{1}$-elliptic projection operator in the jump and average terms for second-order derivatives with an $H^{2}$-elliptic projection operator. Following the idea in [21], we construct an enriching operator to establish the reliability and efficiency of the a posteriori error bound. This enriching operator inherits the beneficial properties of the operators described in [13] for the lowest-order Morley virtual element and is applicable to arbitrary-order cases.
The remainder of the paper is structured as follows. We begin by introducing the continuous variational problem and presenting some useful results in VEM analysis in Section 2. Section 3 details the interior penalty virtual element method, including the virtual element space, degrees of freedom, and the discrete bilinear form, with a particular focus on addressing inhomogeneous boundary value conditions not covered in previous references. In Section 4, we construct an enriching operator and present the corresponding error estimates. Section 5 develops the a posteriori analysis for the IPVEM, providing both upper and lower bounds. Meanwhile, we make a comparison with the analysis of Morley-type virtual elements in [13]. An adaptive IPVEM is presented in Section 6 with several benchmark tests performed to illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM.
2 The continuous variational problem
Let $\Omega$ be a bounded polygonal domain of $\mathbb{R}^{2}$ with boundary $\partial\Omega$. For $f\in L^{2}(\Omega)$, we consider the following boundary value problem of the Kirchhoff plate:
$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=g_{D},\quad\dfrac{\partial u}{\partial\bm{n}}=g_{N}&{\rm on}~{}~{}\partial\Omega.\end{array}\right.$ (2.1)
where $\bm{n}=(n_{1},n_{2})$ is the unit outer normal to $\partial\Omega$, $\Delta$ is the standard Laplacian operator, and $g_{D}\in H^{3/2}(\partial\Omega)$ and $g_{N}\in H^{1/2}(\partial\Omega)$ are the Dirichlet data. Throughout the article we assume $u\in H^{k+1}(\Omega)$ with $k\geq 2$
[…]
$\mathbb{M}_{r}(D):=\Big{\{}\Big{(}\frac{\bm{x}-\bm{x}_{D}}{h_{D}}\Big{)}^{\bm{s}},\quad|\bm{s}|\leq r\Big{\}},$
with the generic monomial denoted by $m_{D}$, where $h_{D}$ is the diameter of $D$, $\bm{x}_{D}$ is the centroid of $D$, and $r$ is a non-negative integer. For the multi-index ${\bm{s}}\in{\mathbb{N}^{d}}$, we follow the usual notation
$\bm{x}^{\bm{s}}=x_{1}^{s_{1}}\cdots x_{d}^{s_{d}},\quad|\bm{s}|=s_{1}+\cdots+s_{d}.$
Conventionally, $\mathbb{M}_{r}(D)=\{0\}$ for $r\leq-1$.
Let $\{\mathcal{T}_{h}\}$ be a family of decompositions of $\Omega$ into polygonal elements $\{K\}$; $h_{K}={\rm diam}(K)$ and $h=\max_{K\in\mathcal{T}_{h}}h_{K}$. Let $\mathcal{V}_{h}$ be the union of all vertices and $\mathcal{E}_{h}$ be the set of all the edges in $\mathcal{T}_{h}$. For each element $K\in\mathcal{T}_{h}$, we denote by $\mathcal{V}(K)$ and $\mathcal{E}(K)$ the sets of vertices and edges of $K$, respectively. Moreover, denote all interior vertices (resp. edges) by $\mathcal{V}_{h}^{0}$ (resp. $\mathcal{E}_{h}^{0}$). The union of boundary edges in $\mathcal{T}_{h}$ is denoted by $\mathcal{E}_{h}^{\partial}$.
Let $e\in\mathcal{E}_{h}^{0}$ be an interior edge shared by two neighbouring elements $K^{-}$ and $K^{+}$. The unit normal vector $\bm{n}_{e}$ associated with $e$ is defined as the outward unit normal of $K=K^{-}$ pointing from $K^{-}$ to $K^{+}$. Let $v$ be a scalar function defined on $e$. We introduce the jump and average of $v$ on $e$ by $[v]=v^{-}-v^{+}$ and $\{v\}=\frac{1}{2}(v^{-}+v^{+})$, where $v^{-}$ and $v^{+}$ are the traces of $v$ on $e$ from the interior and exterior of $K$, respectively. On a boundary edge, we define $[v]=v$ and $\{v\}=v$. From the definition of the averages and jumps, by a direct manipulation, one has $[vw]=[v]\{w\}+\{v\}[w]$ for $e\in\mathcal{E}_{h}^{0}$ and $[vw]=vw=\{v\}[w]=[v]\{w\}$ for a boundary edge, leading to
$\sum_{e\in\mathcal{E}_{h}}\int_{e}[vw]\mathrm{d}s=\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}[v]\{w\}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\int_{e}\{v\}[w]\mathrm{d}s,$ (2.2)
which will be used frequently later. Moreover, for any two quantities $a$ and $b$, “$a\lesssim b$” indicates “$a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and “$a\eqsim b$” abbreviates “$a\lesssim b\lesssim a$”.
The variational formulation of (2.1) with homogeneous boundary value conditions reads: Find $u\in V:=H_{0}^{2}(\Omega)$ such that
$a(u,v)=(f,v),\quad v\in V,$ (2.3)
where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$.
To keep the flexibility of meshes, we will work on meshes satisfying the following assumption (cf. [10, 14]):
  1. H.
    For each $K\in\mathcal{T}_{h}$, there exists a “virtual triangulation” $\mathcal{T}_{K}$ of $K$ such that $\mathcal{T}_{K}$ is uniformly shape regular and quasi-uniform. The corresponding mesh size of $\mathcal{T}_{K}$ is proportional to $h_{K}$. Each edge of $K$ is a side of a certain triangle in $\mathcal{T}_{K}$.
As shown in [14], this condition covers the usual geometric assumptions frequently used in the context of VEMs. According to the standard Dupont-Scott theory (cf. [7]), for all $v\in H^{l}(K)$ ($0\leq l\leq k$) there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
$|v-q|_{m,K}\lesssim h_{K}^{l-m}|v|_{l,K},\quad m\leq l.$ (2.4)
The following trace inequality [7] and the Sobolev inequality [6] are very useful for our forthcoming analysis.
Lemma 2.1 (Trace inequality).
For any $K\in\mathcal{T}_{h}$, it holds
$\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K).$ (2.5)
Lemma 2.2 (Sobolev inequality).
For any $K\in\mathcal{T}_{h}$, it holds
$\|v\|_{\infty,K}\lesssim h_{K}|v|_{2,K}+|v|_{1,K}+h_{K}^{-1}\|v\|_{0,K},\quad v\in H^{2}(K).$ (2.6)
3 The interior penalty virtual element method
This section reviews the interior penalty virtual element space developed in [29] and details the treatment of inhomogeneous boundary value conditions for the IPVEM.
3.1   The IP virtual element spaces
In the construction, the authors in [29] first introduced a $C^{1}$-conforming virtual element space
$\widetilde{V}_{k+2}(K)=\left\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K\right\},\quad k\geq 2.$
This local space can be equipped with the following degrees of freedom (DoFs) (cf. [11, 17]):
  • 1.
    $\tilde{\bm{\chi}}^{p}:$ the values of $v$ at the vertices of $K$,
    $\tilde{\chi}_{z}^{p}(v)=v(z),\quad\mbox{$z$ is a vertex of $K$}.$ (3.1)
  • 2.
    $\tilde{\bm{\chi}}^{g}$ : the values of $h_{z}\nabla v$ at the vertices of $K$,
    $\tilde{\chi}_{z}^{g}(v)=h_{z}\nabla v(z),\quad\mbox{$z$ is a vertex of $K$},$ (3.2)
    where $h_{z}$ is a characteristic length attached to each vertex $z$, for instance, the average of the diameters of the elements having $z$ as a vertex.
  • 3.
    $\tilde{\bm{\chi}}^{e}$ : the moments of $v$ on edges up to degree $k-2$,
    $\tilde{\chi}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-2}(e),\quad e\subset\partial K.$ (3.3)
  • 4.
    $\tilde{\bm{\chi}}^{n}$ : the moments of $\partial_{\bm{n}_{e}}v$ on edges up to degree $k-1$,
    $\tilde{\chi}_{e}^{n}(v)=(m_{e},\partial_{\bm{n}_{e}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-1}(e),\quad e\subset\partial K.$ (3.4)
  • 5.
    $\tilde{\bm{\chi}}^{K}$ : the moments on element $K$ up to degree $k$,
    $\tilde{\chi}_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k}(K).$ (3.5)
Given $v_{h}\in\widetilde{V}_{k+2}(K)$, the usual definition of the $H^{1}$-elliptic projection $\Pi_{K}^{\nabla}v_{h}\in$ $\mathbb{P}_{k}(K)$ is described by the following equations:
$(\nabla\Pi_{K}^{\nabla}v_{h},\nabla q)_{K}&=(\nabla v_{h},\nabla q)_{K},\quad q\in\mathbb{P}_{k}(K),\\ \sum\limits_{z\in\mathcal{V}_{K}}\Pi_{K}^{\nabla}v_{h}(z)&=\sum\limits_{z\in\mathcal{V}_{K}}v_{h}(z),\end{aligned}\right.$ (3.6)
where $\mathcal{V}_{K}$ is the set of the vertices of $K$. By checking the right-hand side of the integration by parts formula:
$(\nabla v_{h},\nabla q)_{K}=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}\int_{e}v_{h}\partial_{\bm{n}_{e}}q\,\mathrm{d}s.$
However, the goal of the IPVEM is to make $\Pi_{K}^{\nabla}v_{h}$ computable by only using the DoFs of $H^{1}$-conforming virtual element spaces given by (cf. [3, 1])
$V_{h}^{1,c}(K):=\{v\in H^{1}(K):\Delta v|_{K}\in\mathbb{P}_{k-2}(K)~{}~{}{\text{in}}~{}~{}K,\quad v|_{\partial K}\in\mathbb{B}_{k}(\partial K)\},$
where
$\mathbb{B}_{k}(\partial K):=\{v\in C(\partial K):v|_{e}\in\mathbb{P}_{k}(e),\quad e\subset\partial K\},$
and the corresponding global virtual element space is
$V_{h}^{1,c}=\{v\in C(\overline{\Omega}):v|_{K}\in V_{h}^{1,c}(K),\,K\in\mathcal{T}_{h}\}\cap H_{0}^{1}(\Omega).$
To do so, Ref. [29] considered the approximation of the right-hand side by some numerical formula. In view of the $\mathbb{P}_{k}$ accuracy, namely $\Pi_{K}^{\nabla}v_{h}=v_{h}$ for $v_{h}\in\mathbb{P}_{k}(K)$, the modified $H^{1}$-projection is defined as
$(\nabla\Pi_{K}^{\nabla}v_{h},\nabla q)_{K}&=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}Q_{2k-1}^{e}(v_{h}\partial_{\bm{n}_{e}}q),\quad q\in\mathbb{P}_{k}(K),\\ \sum_{z\in\mathcal{V}_{K}}\Pi_{K}^{\nabla}v_{h}(z)&=\sum_{z\in\mathcal{V}_{K}}v_{h}(z),\end{aligned}\right.$
with
$Q_{2k-1}^{e}v:=|e|\sum\limits_{i=0}^{k}\omega_{i}v(\bm{x}_{i}^{e})\approx\int_{e}v(s)\mathrm{d}s,$
where $(\omega_{i},\bm{x}_{i}^{e})$ are the $(k+1)$ Gauss-Lobatto quadrature weights and points with $\bm{x}_{0}^{e}$ and $\bm{x}_{k}^{e}$ being the endpoints of $e$.
With the help of $\Pi_{K}^{\nabla}$, we can employ the standard enhancement technique by substituting the redundant moments of $v$ with those of $\Pi_{K}^{\nabla}v$ [1]. The local interior penalty space is defined as
$V_{k}(K)=\{v\in W_{k}(K):\tilde{\chi}^{g}(v)=\tilde{\chi}^{g}(\Pi_{K}^{\nabla}v),\tilde{\chi}^{n}(v)=\tilde{\chi}^{n}(\Pi_{K}^{\nabla}v)\},$
where
$W_{k}(K):=\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k-2}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K,$
$\hskip 113.81102pt(v,q)_{K}=(\Pi_{K}^{\nabla}v,q)_{K},\quad q\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K)\}.$
The associated DoFs are given by
  • 1.
    $\bm{\chi}^{p}:$ the values of $v(z)$, $z\in\mathcal{V}_{K}$;
  • 2.
    $\bm{\chi}^{e}:$ the values of $v(\bm{x}_{i}^{e})$, $i=1,2\cdots,k-1$, $e\subset\partial K$;
  • 3.
    $\bm{\chi}^{K}$ : the moments $|K|^{-1}(m_{K},v)_{K}$, $m_{K}\in\mathbb{M}_{k-2}(K)$.
Furthermore, we use $V_{h}$ to denote the global space of nonconforming virtual element by
[…]
As usual, we can define the $H^{2}$-projection operator $\Pi_{K}^{\Delta}:V_{k}(K)\to\mathbb{P}_{k}(K)$ by finding the solution $\Pi_{K}^{\Delta}v\in\mathbb{P}_{k}(K)$ of
$\begin{cases}a^{K}(\Pi_{K}^{\Delta}v,q)=a^{K}(v,q),\quad q\in\mathbb{P}_{k}(K),\\ \widehat{\Pi_{K}^{\Delta}v}=\widehat{v},\quad\widehat{\nabla\Pi_{K}^{\Delta}v}=\widehat{\nabla v}\end{cases}$ (3.7)
for any given $v\in V_{k}(K)$, where the quasi-average $\widehat{v}$ is defined by
$\widehat{v}=\frac{1}{|\partial K|}\int_{\partial K}v\mathrm{d}s.$
Choosing $q=\Pi_{K}^{\Delta}v$ in (3.7), we can easily observe
$|\Pi_{K}^{\Delta}v|_{2,K}\lesssim|v|_{2,K},\qquad v\in H^{2}(K).$ (3.8)
3.2   The IPVEM with inhomogeneous boundary value conditions
Let $v\in H^{2}(K)$. The integration by parts gives
$(f,v)_{K}=(\Delta^{2}u,v)_{K}=(\nabla^{2}u,\nabla^{2}v)_{K}+\int_{\partial K}\Big{(}\frac{\partial\Delta u}{\partial\bm{n}_{K}}v+\frac{\partial^{2}u}{\partial\bm{n}_{K}\partial\bm{t}_{K}}\frac{\partial v}{\partial\bm{t}_{K}}\Big{)}\mathrm{d}s-\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v}{\partial\bm{n}_{K}}\mathrm{d}s.$
If $v$ is $C^{0}$-continuous, the second term on the right-hand side naturally vanishes when we sum the equation over all polygons in $\mathcal{T}_{h}$. However, when $v=v_{h}$ is a function in the interior penalty virtual element space $V_{h}$, which is not continuous, the second term contributes to the summation. The IPVEM in [29] neglects this contribution as it introduces a consistency error with a desired convergence order as shown in [29, 18], which gives
$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s.$
Using the formula in (2.2) and noting that $u\in H^{3}(\Omega)$, we get
$\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s$ $=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}\Big{[}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\{}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{\}}\mathrm{d}s$
$=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
yielding
$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
To ensure both the symmetry and $C^{1}$-continuity of the discrete variational problem, we rewrite it as
$(f,v_{h})-\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
$\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
where we have used the fact that $[u]_{e}=0$ and $[\frac{\partial u}{\partial\bm{n}_{e}}]_{e}=0$ for $e\in\mathcal{E}_{h}^{0}$.
Following the standard procedure in the VEMs, we define the discrete bilinear form as
$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$
with
$S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w)=h_{K}^{-2}\sum\limits_{i=1}^{n_{K}}\chi_{i}(v-\Pi_{K}^{\Delta}v)\chi_{i}(w-\Pi_{K}^{\Delta}w),$
where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. Let
$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[v][w]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$ (3.9)
with $\lambda_{e}\geq 1$ being some edge-dependent parameter, and the additional two terms are given by
$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
Here and bellows, we define the piecewise $H^{2}$ projector $\Pi_{h}^{\Delta}$ by setting $\Pi_{h}^{\Delta}|_{K}=\Pi_{K}^{\Delta}$ for all $K\in\mathcal{T}_{h}$.
The IPVEM for (2.1) with inhomogeneous boundary value conditioins can be described as follows: Find $u_{h}\in V_{h}^{g}=\{v\in V_{h}:v=I_{h}g_{D}~{}{\rm on}~{}\partial\Omega\}$ such that
$\mathcal{A}_{h}(u_{h},v_{h})=F_{h}(v_{h}),\quad v_{h}\in V_{h},$ (3.10)
where $g_{D}$ has been extended to the whole domain. The bilinear form is
$\mathcal{A}_{h}(u_{h},v_{h})=a_{h}(u_{h},v_{h})+J_{1}(u_{h},v_{h})+J_{2}(u_{h},v_{h})+J_{3}(u_{h},v_{h}),$ (3.11)
where $a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w)$. The right-hand side is
$F_{h}(v_{h})=\langle f_{h},v_{h}\rangle+\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{(}-\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}+\frac{\lambda_{e}^{3}}{|e|^{3}}[u][v_{h}]+\frac{\lambda_{e}}{|e|}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{)}\mathrm{d}s,$
where
$\langle f_{h},v_{h}\rangle:=\sum\limits_{K\in\mathcal{T}_{h}}\int_{K}f\,\Pi_{0,K}^{k}v_{h}\mathrm{d}x=\sum\limits_{K\in\mathcal{T}_{h}}\int_{K}\Pi_{0,K}^{k}f\,v_{h}\mathrm{d}x,$ (3.12)
with $\Pi_{0,K}^{k}$ being the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. Then for $f\in H^{k-2}(\Omega)$ with $k\geq 2$, we have
$\langle f-f_{h},v_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}(f,v_{h}-\Pi_{0,K}^{k}v_{h})=\sum_{K\in\mathcal{T}_{h}}(f-\Pi_{0,K}^{k}f,v_{h}-\Pi_{0,K}^{k}v_{h})$
$\lesssim h^{k-2}\|f\|_{k-2}h^{2}|v_{h}|_{2,h}\lesssim h^{k}\|f\|_{k-2}|v_{h}|_{2,h},$
which means
$\|f-f_{h}\|_{V^{\prime}_{h}}\lesssim h^{k}\|f\|_{k-2}.$ (3.13)
In what follows, we define
$\|w\|_{h}^{2}:=|w|_{2,h}^{2}+J_{1}(w,w).$ (3.14)
With the mesh-dependent parameter $\lambda_{e}$ chosen as in [18], we can deduce the stability result by using the same arguments. We omit the details with the results described as follows.
    $k$-consistency: for all $v\in V_{k}(K)$ and $q\in\mathbb{P}_{k}(K)$, it holds that
    $a_{h}^{K}(v,q)=a^{K}(v,q).$ (3.15)
    Stability: there exist two positive constants $\alpha_{*}$ and $\alpha^{*}$, independent of $h$, such that
    $\alpha_{*}a^{K}(v,v)\leq a_{h}^{K}(v,v)\leq\alpha^{*}a^{K}(v,v)$ (3.16)
    for all $v\in V_{k}(K)$.
For later use, we provide the following norm equivalence with the proof given in [29].
Lemma 3.1.
For $m=0,1,2$, there holds
$|v-\Pi_{K}^{\Delta}v|_{m,K}\eqsim h_{K}^{1-m}\|\bm{\chi}(v-\Pi_{K}^{\Delta}v)\|_{\ell^{2}},\qquad v\in V_{k}(K).$ (3.17)
4 Enriching operator
This section focuses on proposing and analyzing an enriching operator $E_{h}$ from $V_{h}$ to $H^{2}$-conforming virtual element spaces. The enriching operator is a special kind of quasi-interpolation operator, which connects a finite dimensional space to another one with higher order regularity. We refer to [5, 9, 19, 22, 20, 8] and the references therein for the construction and application of such operators in finite element methods. Such an operator in the virtual element context was first proposed in [21] and applied in the medius error analysis of some nonconforming VEMs. This idea was also adopted in [16] for the a posteriori error estimation for a $C^{1}$ virtual element method of Kirchhoff plates.
Let us recall the $H^{2}$-conforming virtual element space $V_{k}^{2,c}(K)$ introduced in [11, 17]. For $k\geq 3$, define
${V_{k}^{2,c}}(K)=\{v\in H^{2}(K):\Delta^{2}v|_{K}\in\mathbb{P}_{k-4}(K),~{}v|_{e}\in\mathbb{P}_{k}(e),~{}\partial_{\bm{n}}v|_{e}\in{\mathbb{P}_{k-1}}(e),~{}e\subset\partial K\},$
while for the lowest order $k=2$, the space is modified as
$V_{2}^{2,c}(K)=\Big{\{}v\in H^{2}(K):\Delta^{2}v|_{K}=0,~{}v|_{e}\in\mathbb{P}_{3}(e),~{}\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{1}(e),~{}e\subset\partial K\Big{\}}.$
The DoFs are:
  • 1.
    The values of $v(z)$ at the vertices of $K$.
  • 2.
    The values of $h_{z}\partial_{1}v(z)$ and $h_{z}\partial_{2}v(z)$ at the vertices of $K$, where $h_{z}$ is a characteristic length attached to each vertex $z$, for instance, the average of the diameters of the elements having $z$ as a vertex.
  • 3.
    The moments of $v$ on edges up to degree $k-4$,
    $\chi_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e).$
  • 4.
    The moments of $\partial_{\bm{n}}v$ on edges up to degree $k-3$,
    $\chi_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e).$
  • 5.
    The moments on element $K$ up to degree $k-4$,
    $\chi_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k-4}(K).$
Note that the above moments vanish in the lowest order case $k=2$.
The global virtual element spaces are then defined as
$V_{h}^{2,c}:=\Big{\{}v\in H_{0}^{2}(\Omega)\cap C^{1}(\overline{\Omega}):v|_{K}\in V_{k}^{2,c}(K),~{}~{}K\in\mathcal{T}_{h}\Big{\}}.$
Let ${\mathcal{E}_{a}^{0}}$ and $\mathcal{E}_{e}^{0}$ denote the interior vertices and edges of $\mathcal{T}_{h}$, respectively. The global dual space is defined as
$\mathcal{N}=\text{span}\Big{\{}\mathcal{N}_{a},\mathcal{N}_{\nabla},\mathcal{N}_{e}^{k-4},\mathcal{N}_{n_{e}}^{k-3},\mathcal{N}_{K}^{k-4}\Big{\}}$
with global DoFs given by
  • 1.
    $\mathcal{N}_{a}$: the values at the interior vertices of $\mathcal{T}_{h}$,
    $\mathcal{N}_{a}^{z}(v)=v(z),\quad z\in\mathcal{E}_{a}^{0}.$
  • 2.
    $\mathcal{N}_{\nabla}$: the gradient values at interior vertices of $\mathcal{T}_{h}$,
    $\mathcal{N}_{\nabla}^{z}(v)=h_{z}\nabla v(z),\quad z\in\mathcal{E}_{a}^{0}.$
  • 3.
    ${\mathcal{N}_{e}^{k-4}}$: the moments of $v$ on interior edges up to degree $k-4$,
    $\mathcal{N}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e),\quad e\in\mathcal{E}_{e}^{0}.$
  • 4.
    $\mathcal{N}_{n_{e}}^{k-3}$: the moments of $\partial_{\bm{n}}v$ on interior edges up to degree $k-3$,
    $\mathcal{N}_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e),\quad e\in\mathcal{E}_{e}^{0}.$
  • 5.
    $\mathcal{N}_{K}^{k-4}$: the moments on element $K$ up to degree $k-4$,
    $\mathcal{N}_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k-4}(K).$
Relabel the d.o.f.s by a single index $i=1,2,\ldots,N^{2,c}:=\dim V_{h}^{2,c}$, which are associated with a set of shape basis functions $\{\Phi_{j}\}$ of $V_{h}^{2,c}$ such that $\mathcal{N}_{i}(\Phi_{j})=\delta_{ij}$ for $i,j=1,\cdots,N^{2,c}$. Then every function $v\in V_{h}^{2,c}$ can be expressed as
$v(x)=\sum_{i=1}^{N^{2,c}}\mathcal{N}_{i}(v)\Phi_{i}(x).$ (4.1)
For a vertex $z$ of $\mathcal{T}_{h}$, let $\omega(z)$ denote the union of all elements in $\mathcal{T}_{h}$ sharing the point $z$. For an edge $e$ in $\mathcal{T}_{h}$, let $\omega(e)$ be the union of all elements in $\mathcal{T}_{h}$ sharing the edge $e$. Let $N(z)$ and $N(e)$ denote the number of elements in $\omega(z)$ and $\omega(e)$, respectively.
For every nonconforming VEM function $\phi\in V_{h}$, we shall construct an associated conforming counterpart $E_{h}\phi\in V_{h}^{2,c}\cap H_{0}^{2}(\Omega)$ by using the global expansion (4.1), i.e.,
$(E_{h}\phi)(x)=\sum\limits_{i=1}^{N^{2,c}}\mathcal{N}_{i}(E_{h}\phi)\Phi_{i}(x).$
Here, the values of the DoFs are determined as follows:
  1. 1.
    For the values at interior vertices,
    $\mathcal{N}_{a}^{z}(E_{h}\phi)=(E_{h}\phi)(z):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
  2. 2.
    For the gradient values at interior vertices,
    $\mathcal{N}_{\nabla}^{z}(E_{h}\phi):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}h_{z}\nabla\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
  3. 3.
    For the moments of $v$ on interior edges,
    $\mathcal{N}_{e}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}|e|^{-1}\int_{e}m_{e}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}\mathrm{d}s,\quad m_{e}\in\mathbb{M}_{k-4}(e).$
  4. 4.
    For the moments of ${\partial_{\bm{n}}}v$ on interior edges,
    $\mathcal{N}_{n_{e}}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}\int_{e}m_{e}\partial_{\bm{n}}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}{\text{d}}s,\quad m_{e}\in\mathbb{M}_{k-3}(e).$
  5. 5.
    For the moments on element $K$,
    $\mathcal{N}_{K}(E_{h}\phi):=|K|^{-1}\int_{K}m_{K}\Pi_{h}^{\Delta}\phi|_{K}{\mathrm{d}}x,\quad m_{K}\in\mathbb{M}_{k-4}(K).$
Lemma 4.1.
For any $v\in V_{h}$ and $E_{h}v$ defined above, there holds
$|v-E_{h}v|_{2,h}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)},$ (4.2)
Proof.
According to the triangle inequality
$|v-E_{h}v|_{2,h}\leq|v-\Pi_{h}^{\Delta}v|_{2,h}+|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}$
and the norm equivalence (3.17), it reduces to bound the second term on the right-hand side. Since $\Pi_{h}^{\Delta}v-E_{h}v\in V_{k}^{2,c}(K)$, by the inverse inequality and norm equivalence in [21], one gets
$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,K}\lesssim h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\lesssim h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}},\qquad K\in\mathcal{T}_{h}.$
where $\bm{\chi}^{c}$ is the set of DoFs on $K$ for the $C^{1}$-continuous space. In the following, we only provide the details of bounding the values and its gradient values of the function $\Pi_{h}^{\Delta}v-E_{h}v$ at vertices since the argument implies the treatment for moments.
(1) Let $z\in\mathcal{V}_{h}^{0}$ be an interior vertex. Assume that $K_{1}=K,K_{2},\cdots,K_{L}$ share the node $z$, and denote $(\Pi_{h}^{\Delta}\phi)_{i}=\Pi_{h}^{\Delta}\phi|_{K_{i}}$, where $K_{i}$ and $K_{i+1}$ are two neighboring elements. We have
$(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
$=(\Pi_{h}^{\Delta}v)_{1}(z)-\frac{1}{L}((\Pi_{h}^{\Delta}v)_{1}+\cdots+(\Pi_{h}^{\Delta}v)_{L})(z)$
$=\frac{1}{L}(((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2})+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{3})+\cdots+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{L}))(z).$
Since $L$ is uniformly bounded and
$(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{j}(z)=\sum\limits_{i=1}^{j-1}((\Pi_{h}^{\Delta}v)_{i}-(\Pi_{h}^{\Delta}v)_{i+1})(z),$
it suffices to consider the term $(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)$. Apply the inverse inequality of polynomials to get
$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|\leq\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{\infty,e}\lesssim h_{e}^{-1/2}\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}.$ (4.3)
Noting that $v$ is not continuous across $e$, we denote by $v_{i}$ the restriction of $v$ on $K_{i}$. Then one has
$\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$ $=\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}+v_{1}-v_{2}+v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$
$\leq\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}\|_{0,e}+\|v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}+\|[v]\|_{0,e}.$
By the trace inequality (2.5),
$\|v-\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1/2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
If $z$ is a vertex on the domain boundary, then $E_{h}(v)(z)=0=v(z)$, which gives
$|(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|=|\Pi_{h}^{\Delta}v(z)-v(z)|\leq\|\Pi_{h}^{\Delta}v-v\|_{\infty,K}.$
By the Sobolev inequality (2.6),
$\|v-\Pi_{h}^{\Delta}v\|_{\infty,K}\leq h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
$\Big{(}h_{K}^{-1}(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2},$
where $e_{i}$ is the edge sharing by $K_{i}$ and $K_{i+1}$.
(2) For the gradient value at an interior vertex $z\in\mathcal{V}^{0}$, one has
$h_{z}\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
$=$ $h_{z}(\nabla\Pi_{h}^{\Delta}v)_{1}(z)-\frac{h_{z}}{L}((\nabla\Pi_{h}^{\Delta}v)_{1}+\cdots+(\nabla\Pi_{h}^{\Delta}v)_{L})(z)$
$=$ $\frac{h_{z}}{L}(((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{2})+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{3})+\cdots+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{L}))(z)$
and we only need to consider the term $h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$. By the inverse inequality of polynomials,
$h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$ $\lesssim h_{z}h_{e}^{-1/2}|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}$
$\lesssim h_{e}^{1/2}\Big{(}|(\Pi_{h}^{\Delta}v)_{1}-v_{1}|_{1,e}+|v_{2}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}+|[v]|_{1,e}\Big{)}.$
Using the trace inequality, we get
$|v-\Pi_{h}^{\Delta}v|_{1,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{2,K}+h_{K}^{-1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}.$
For the jump term $|[v]|_{1,e}$, noting that $[\partial_{\bm{t}_{e}}v]$ is a polynomial on $e$, we obtain from the inverse inequality for polynomials to get $\|[\partial_{\bm{t}_{e}}v]\|_{e}\lesssim h_{e}^{-1}\|[v]\|_{e}$ and hence
$|[v]|_{1,e}\lesssim\|[\partial_{\bm{n}_{e}}v]\|_{0,e}+h_{e}^{-1}\|[v]\|_{e}.$
For a boundary vertex $z$, we have $\nabla E_{h}(v)(z)=\nabla v(z)=0$. Let $e$ be the associated boundary edge. By the inverse inequality for polynomials and the trace inequality,
$h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|$ $=h_{z}|\nabla\Pi_{h}^{\Delta}v(z)|\leq h_{K}\|\nabla\Pi_{h}^{\Delta}v\|_{\infty,e}\lesssim h_{e}^{1/2}\|\nabla\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v\|_{0,e}$
$=h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v-v\|_{0,e}\lesssim|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
$\Big{(}h_{K}^{-1}h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\Big{(}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2}+\frac{1}{|e_{i}|}\|[\partial_{\bm{n}_{e_{i}}}v]\|_{0,e_{i}}^{2}\Big{)}.$
This completes the proof. ∎
Remark 4.1.
Our result aligns with the finite element methods since the first term on the right-hand side of (4.2) vanishes in this case (see [9, 8, 12] for instance).
For the medius error analysis, we should bound the (semi-)norm of $v-E_{h}v$ by $|v|_{2,h}$. Here, we present the result for $\|v-E_{h}v\|_{0,K}$ as an example, even though this estimate will not be needed in the subsequent analysis.
Lemma 4.2.
For any $v\in V_{h}$ and $E_{h}v$ defined above, there holds
$\sum\limits_{K\in\mathcal{T}_{h}}h_{K}^{-4}\|v-E_{h}v\|_{0,K}^{2}\lesssim|v|_{2,h}^{2}.$
Proof.
From the triangle inequality
$\|v-E_{h}v\|_{0,K}\leq\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$
and the standard estimate
$h_{K}^{-2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\lesssim|v|_{2,K},$
it remains to bound the term $\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$. By the norm equivalence,
$h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\simeq h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}}.$
Let $z$ be an interior vertex. As in the proof of Lemma 4.1, one only needs to consider $|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$. Noting that $v$ is continuous at $z$ (i.e., $v_{1}(z)=v_{2}(z)$), by applying the Sobolev inequality (2.6), we obtain
$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$ $\leq|(\Pi_{h}^{\Delta}v)_{1}(z)-v_{1}(z)|+|v_{2}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$
$\lesssim\sum_{K=K_{1},K_{2}}\Big{(}h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\Big{)}.$
The proof is finished by using the error estimate for $\Pi_{h}^{\Delta}$. ∎
5 A posteriori error estimation
For the a posteriori error estimation, without loss of generality, we simply assume homogeneous boundary value conditions. By reviewing the calculation in Section 5.1 of [16], we can incorporate the computable components of the local error estimator $\eta_{K}$ as follows:
$\eta_{0,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{0,e}^{2}\Big{)}^{1/2},\qquad\eta_{0,e}=\frac{\lambda_{e}^{3/2}}{|e|^{3/2}}\|[u_{h}]\|_{0,e},$
$\eta_{1,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2},\qquad\eta_{1,e}=\frac{\lambda_{e}^{1/2}}{|e|^{1/2}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e},$
$\eta_{2,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{2,e}^{2}\Big{)}^{1/2},\qquad\eta_{2,e}=|e|^{1/2}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{3,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2},\qquad\eta_{3,e}=|e|^{3/2}\Big{\|}\Big{[}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}}+\frac{\partial^{3}(\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
$\eta_{4,K}=\big{(}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})\big{)}^{1/2},$
$\eta_{5,K}:=h_{K}^{2}\|f-f_{h}\|_{0,K},$
$\eta_{6,K}=h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K},$ (5.1)
where $f_{h}|_{K}=\Pi_{0,K}^{k}f$ and $\omega(K)$ denotes the union of all elements in $\mathcal{T}_{h}$ sharing some edges or points with $K$. The first term results from the penalty term which measures the extent to which $u_{h}$ fails to be in $H_{0}^{2}(\Omega)$. The remaining five terms successively indicate the residual of the normal moment jump and the jump in the effective shear force along interior edges, the virtual element consistency residual, the element data oscillation and the element residual.
The local and global error estimators are respectively defined by
$\eta_{K}(u_{h})=\Big{(}\sum\limits_{i=0}^{6}\eta_{i,K}^{2}\Big{)}^{1/2}$
and
$\eta(u_{h})=\eta(u_{h},\mathcal{T}_{h})=:\Big{(}\sum\limits_{K\in\mathcal{T}_{h}}\eta_{K}^{2}(u_{h})\Big{)}^{1/2}.$
5.1   Upper bound
Theorem 5.1.
Let $u\in H_{0}^{2}(\Omega)\cap H^{k+1}(\Omega)$ with $k\geq 2$. Suppose that $u$ and $u_{h}$ are the solutions to (2.1) and (3.10), respectively. Then it holds
$\|u-u_{h}\|_{h}\lesssim\eta(u_{h}).$
Proof.
Let $e_{h}=u-u_{h}$ and $E_{h}$ be the enriching operator. Using the triangle inequality and (4.2), one has
$\|e_{h}\|_{h}$ $\leq|u-E_{h}u_{h}|_{2}+|E_{h}u_{h}-u_{h}|_{2,h}+J_{1}(e_{h},e_{h})^{1/2}$ (5.2)
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})$
$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|^{3}}\|[u_{h}]\|_{0,e}^{2}+\sum\limits_{e\in\mathcal{E}_{h}}\frac{1}{h_{e}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}+J_{1}(e_{h},e_{h})^{1/2}$
$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}.$
Therefore, we only need to consider the first term $|u-E_{h}u_{h}|_{2}$ in the last equation.
Let $\phi=u-E_{h}u_{h}$ and denote by $I_{h}:H^{2}(\Omega)\to V_{h}$ the interpolation operator. According to the definitions of the continuous and discrete variational problems, one gets
$|u-E_{h}u_{h}|_{2}^{2}=a(u-E_{h}u_{h},\phi)$
$=\sum_{K\in\mathcal{T}_{h}}\Big{(}a^{K}(u_{h}-E_{h}u_{h},\phi)-a^{K}(u_{h},\phi-I_{h}\phi)\Big{)}+a(u,\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h},I_{h}\phi)$
$=\sum_{K\in\mathcal{T}_{h}}\Big{(}a^{K}(u_{h}-E_{h}u_{h},\phi)-a^{K}(u_{h},\phi-I_{h}\phi)\Big{)}+(f,\phi)-(f_{h},I_{h}\phi)$
$\qquad+a_{h}(u_{h},I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h},I_{h}\phi)+J_{1}(u_{h},I_{h}\phi)+J_{2}(u_{h},I_{h}\phi)+J_{3}(u_{h},I_{h}\phi).$
The definition of $\Pi_{h}^{\Delta}$ gives
$a_{h}(u_{h},I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h},I_{h}\phi)$
$=$ $\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\Delta}u_{h},\Pi_{h}^{\Delta}I_{h}\phi)+\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},I_{h}\phi-\Pi_{h}^{\Delta}I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h},I_{h}\phi)$
$=$ $\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\Delta}u_{h}-u_{h},I_{h}\phi)+\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},I_{h}\phi-\Pi_{h}^{\Delta}I_{h}\phi)$
Therefore, we can decompose $|u-E_{h}u_{h}|_{2,h}^{2}$ as
$a(u-E_{h}u_{h},\phi)=I_{1}+I_{2}+I_{3}+I_{4}+J_{1}(u_{h},I_{h}\phi)+J_{2}(u_{h},I_{h}\phi)+J_{3}(u_{h},I_{h}\phi),$ (5.3)
where
$I_{1}=\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h}-E_{h}u_{h},\phi),$
$I_{2}=(f_{h},\phi-I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)$ (5.4)
$I_{3}=-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\phi)+\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},I_{h}\phi-\Pi_{h}^{\Delta}I_{h}\phi)$
$I_{4}=(f-f_{h},\phi).$
For the term $I_{1}$, using the Cauchy-Schwarz inequality and the estimate (4.2), one gets
$I_{1}$ $\leq\sum_{K\in\mathcal{T}_{h}}|u_{h}-E_{h}u_{h}|_{2,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-E_{h}u_{h}|_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}$
$\lesssim\Big{(}\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
For the term $I_{2}$, using the integration by parts we can split it into four terms as
$I_{2}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}},\phi-I_{h}\phi\Big{)}_{\partial K}$
$\qquad-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$\qquad+\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)$
$=:I_{2,1}+I_{2,2}+I_{2,3}+I_{2,4}.$
For $I_{2,1}$, by the interpolation error estimate,
$I_{2,1}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}\leq\sum_{K\in\mathcal{T}_{h}}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}\|\phi-I_{h}\phi\|_{0,K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{6,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For the sake of brevity, we set $Q_{3}(u_{h})=\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}}$ in $I_{2,2}$. From the equality (2.2), one gets
$I_{2,2}=$ $\sum_{K\in\mathcal{T}_{h}}(Q_{3}(u_{h}),\phi-I_{h}\phi)_{\partial K}$
$=$ $\sum\limits_{e\in\mathcal{E}_{h}^{0}}([Q_{3}(u_{h})],\{\phi-I_{h}\phi\})_{e}+\sum\limits_{e\in\mathcal{E}_{h}}(\{Q_{3}(u_{h})\},[\phi-I_{h}\phi])_{e}.$
On the other hand, noting that $I_{h}\phi|_{e}\in\mathbb{P}_{k+2}(e)$ and $Q_{3}(u_{h})|_{e}\in\mathbb{P}_{k-3}(e)$, one has ($(\{Q_{3}(u_{h})\},[I_{h}\phi])_{e}=0$, since $I_{h}\phi$ is continuous at the Gauss-Lobatto points on $e$ and the quadrature formula is exact for polynomials of up to degree $12k-$ 1. Combing with $\phi\in H_{0}^{2}(\Omega)$, we derive
$\sum\limits_{e\in\mathcal{E}_{h}}(\{Q_{3}(u_{h})\},[\phi-I_{h}\phi])_{e}=0.$
The continuity of $I_{h}\phi$ at the Gauss-Lobatto points on $e$ implies that
$(\phi-I_{h}\phi)(z)=0,\quad z\in\mathcal{V}_{h},$
which indicates
$I_{2,4}=\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)=0.$
Hence, it follows from the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates that
$I_{2,2}$ $=\sum\limits_{e\in\mathcal{E}_{h}^{0}}([Q_{3}(u_{h})],\{\phi-I_{h}\phi\})_{e}\leq\sum\limits_{e\in\mathcal{E}_{h}^{0}}\|[Q_{3}(u_{h})]\|_{0,e}\|\{\phi-I_{h}\phi\}\|_{0,e}$
$\lesssim\sum\limits_{e\in\mathcal{E}_{h}^{0}}\|[Q_{3}(u_{h})]\|_{0,e}(h_{e}^{-1/2}\|\{\phi-I_{h}\phi\}\|_{0,\omega(e)}+h_{e}^{1/2}|\{\phi-I_{h}\phi\}|_{1,\omega(e)})$
$\lesssim\sum\limits_{e\in\mathcal{E}_{h}^{0}}h_{e}^{3/2}\|[Q_{3}(u_{h})]\|_{0,e}|\phi|_{2,\omega(e)}\lesssim\Big{(}\sum_{e\in\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{3,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For $I_{2,3}$, noting that $[\frac{\partial\phi}{\partial\bm{n}}]_{e}=0$, from (2.2), we have
$I_{2,3}$ $=-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}\Big{(}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-\sum_{e\in\mathcal{E}_{h}}(\{\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}\},[\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}])_{e}\Big{)}$
$=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-J_{2}(u_{h},I_{h}\phi).$
Similarly, by using the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates, we can get
$I_{2,3}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
Collecting the above estimates to derive
$I_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{2,K}^{2}+\eta_{3,K}^{2}+\eta_{6,K}^{2})\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
For the term $I_{3}$, by the norm equivalence, we obtain
$I_{3}$ $=-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\phi)+\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},I_{h}\phi-\Pi_{h}^{\Delta}I_{h}\phi)$
$\lesssim\sum_{K\in\mathcal{T}_{h}}\Big{(}h_{K}^{-2}\|{\bm{\chi}}(u_{h}-\Pi_{h}^{\Delta}u_{h})\|_{\ell^{2}}\|{\bm{\chi}}({I_{h}}\phi-\Pi_{h}^{\Delta}({I_{h}}\phi))\|_{\ell^{2}}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|\phi|_{2,K}\Big{)}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}\Big{(}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|{I_{h}}\phi-\Pi_{h}^{\Delta}({I_{h}}\phi)|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|\phi|_{2,K}\Big{)}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}\Big{)}^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}\big{(}|\phi|_{2,K}^{2}+|\phi-{I_{h}}\phi|_{2,K}^{2}\big{)}\Big{)}^{1/2}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
For $I_{4}$, it is easy to get
$I_{4}=(f-f_{h},\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{5,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
For $J_{1}(u_{h},I_{h}\phi)$, by the interpolation error estimate, we have
$J_{1}(u_{h},I_{h}\phi)\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{e\in\mathcal{E}_{h}}|e|^{-1}\Big{\|}\Big{[}\frac{\partial(I_{h}\phi-\phi)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
For $J_{3}(u_{h},I_{h}\phi)$, the boundedness of $\Pi_{h}^{\Delta}$ gives
$J_{3}(u_{h},I_{h}\phi)$ $=\Big{|}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}I_{h}\phi}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{|}\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{2,e}^{2})^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\Delta}\phi|_{2,K}^{2}\Big{)}^{1/2}$
$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$ (5.5)
The combination of the above estimates yields
$|u-E_{h}u_{h}|_{2}^{2}=a(u-E_{h}u_{h},\phi)\lesssim\eta(u_{h})|\phi|_{2}=\eta(u_{h})|u-E_{h}u_{h}|_{2},$
as required. ∎
5.2   Lower bound
By means of the technique of bubble functions frequently used in the a posteriori error analysis [21, 24], we can also derive the following lower bound of the a posteriori error estimator.
Theorem 5.2.
Let $u\in V$ be the solution of (2.3). Then, for any $w_{h}\in V_{h}$, $K\in\mathcal{T}_{h}$ and $e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}$, there hold
$h^{4}_{K}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}w_{h}\|_{0,K}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,K}^{2}+\eta_{5,K}^{2},$ (5.6)
$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2},$ (5.7)
$h_{e}^{3}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}}+\frac{\partial^{3}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2}.$ (5.8)
Proof.
We divide the proof into three steps. For brevity, we omit the subscript $e$ in $\bm{t}_{e}$ and $\bm{n}_{e}$. For the sake of brevity, we denote $v=\Pi_{h}^{\Delta}w_{h}$ in what follows.
Step 1: For any $K\in\mathcal{T}_{h}$, assume $\mathcal{T}_{K}$ is the auxiliary triangulation of $K$, and let $b_{K}=\sum_{\tau\in\mathcal{T}_{K}}(\lambda_{1,\tau}\lambda_{2,\tau}\lambda_{3,\tau})^{2}$ be the element bubble function defined on $K$, where $\lambda_{i,\tau}$($i=1,2,3$) are the nodal basis functions at vertices of virtual triangles $\tau\in\mathcal{T}_{K}$. Set the polynomial $\phi|_{K}=b_{K}(f_{h}-\Delta^{2}v)|_{K}$ and extend it to be zero in $\Omega\backslash{K}$. Denote the auxiliary function as $\phi$, and it is obviously that $\phi\in H^{2}_{0}(\Omega)$ and $\partial_{\bm{n}}\phi|_{\partial{K}}=\phi|_{\partial{K}}=0$, which yields
$(f-\Delta^{2}v,\,\phi)$ $=a(u,\,\phi)-(\Delta^{2}v,\,\phi)_{K}=a^{K}(u-v,\phi).$
Then, using the inverse inequalities of polynomials and the technique of bubble functions, there holds
$\|f_{h}-\Delta^{2}v\|_{0,K}^{2}$ $\lesssim(f_{h}-\Delta^{2}v,\,\phi)_{K}=(f_{h}-f,\,\phi)_{K}+(f-\Delta^{2}v,\,\phi)_{K}$
$=(f_{h}-f,\,\phi)_{K}+a^{K}(u-v,\,\phi)$
$\leq\|f-f_{h}\|_{0,K}\|\phi\|_{0,K}+|u-v|_{2,K}|\phi|_{2,K}$
$\lesssim(\|f-f_{h}\|_{0,K}+h_{K}^{-2}|u-v|_{2,K})\|\phi\|_{0,K}$
$\lesssim(\|f-f_{h}\|_{0,K}+h_{K}^{-2}|u-v|_{2,K})\|f_{h}-\Delta^{2}v\|_{0,K}.$
The required estimate (5.6) is obtained easily.
Step 2. For any interior edge $e\in\mathcal{E}_{h}^{0}$ shared by two elements $K^{+}$ and $K^{-}$, assume that $\tau^{\pm}$ are two triangles sharing the edge $e$ in the auxiliary triangulation $\mathcal{T}_{K^{+}}\cup\mathcal{T}_{K^{-}}$, viz. $e=\tau^{+}\cap\tau^{-}$, where $\tau^{+}\in\mathcal{T}_{K^{+}}$ and $\tau^{-}\in\mathcal{T}_{K^{-}}$. Denote $T=\tau^{+}\cup\tau^{-}$. Let $\lambda_{1,e}^{\pm}$ and $\lambda_{2,e}^{\pm}$ be the corresponding nodal basis functions of two endpoints of $e$ in $\tau^{\pm}$, and denote the unit outward normal vector of $\tau^{+}$ and $\tau^{-}$ along $e$ by $\bm{n}^{+}$ and $\bm{n}^{-}$, respectively.
According to the technique of bubble functions, we need to construct an auxiliary polynomial $\phi\in H^{2}_{0}(T)$ satisfying $\partial_{\bm{n}}\phi|_{\partial{T}\setminus{e}}=0$ and
$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{e}$ (5.9)
Since the jump $[\frac{\partial^{2}v}{\partial\bm{n}^{2}}]|_{e}$ is a polynomial on $e$, we extend it to $T$, denoted by $E_{T}([\frac{\partial^{2}v}{\partial\bm{n}^{2}}])$, so that this function is constant along the lines perpendicular to $e$. Let $\phi=b_{e}E_{T}([\frac{\partial^{2}v}{\partial\bm{n}^{2}}])l_{e}$, where $l_{e}$ is the function of line $e$, and $b_{e}$ is edge bubble function of $e$ defined by $b_{e}=(\lambda_{1,e}^{+}\lambda_{2,e}^{+}\lambda_{1,e}^{-}\lambda_{2,e}^{-})^{2}$. A simple manipulation yields $\partial_{\bm{n}}\phi|_{e}=b_{e}[\frac{\partial^{2}v}{\partial\bm{n}^{2}}]|_{e}$ and $\phi|_{e}=0$. Next, by (5.9), there holds
$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}=a^{T}(v,\,\phi)-(\Delta^{2}v,\,\phi)_{T}.$ (5.10)
Since $\phi\in H^{2}_{0}(T)$ can be viewed as a function in $H_{0}^{2}(\Omega)$ by zero extension, we have
$a^{T}(u,\,\phi)=(f,\phi)_{T}.$ (5.11)
Using (5.10), (5.11), the Cauchy-Schwarz inequality and the inverse inequalities of polynomials, we find
$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}$ $\lesssim a^{T}(v-u,\,\phi)+(f-\Delta^{2}v,\,\phi)_{T}$
$\lesssim\sum_{\tau=\tau^{+},\tau^{-}}\Big{(}|v-u|_{2,\tau}|\phi|_{2,\tau}+\|f-\Delta^{2}v\|_{0,\tau}\|\phi\|_{0,\tau}\Big{)}$
$\lesssim\sum_{\tau=\tau^{+},\tau^{-}}\Big{(}h_{\tau}^{-2}|v-u|_{2,\tau}+\|f-\Delta^{2}v\|_{0,\tau}\Big{)}\|\phi\|_{0,\tau}.$ (5.12)
On the other hand, the inverse inequalities of polynomials imply
$\|\phi\|_{0,\tau^{\pm}}$ $=h_{\tau^{\pm}}\Big{\|}b_{e}E_{T}\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{|}_{e}\Big{)}\frac{l_{e}}{h_{\tau^{\pm}}}\Big{\|}_{0,\tau^{\pm}}\leq h_{\tau^{\pm}}\Big{\|}E_{T}\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{|}_{e}\Big{)}\Big{\|}_{0,\tau^{\pm}}$
$\lesssim h_{\tau^{\pm}}^{2}\Big{\|}E_{T}\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{|}_{e}\Big{)}\Big{\|}_{\infty,\tau^{\pm}}\lesssim h_{\tau^{\pm}}^{2}\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{\infty,e}$
$\lesssim h_{\tau^{\pm}}^{3/2}\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{0,e}\lesssim h_{K^{\pm}}^{3/2}\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{0,e}.$ (5.13)
Hence,
$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim\sum_{\tau=\tau^{+},\tau^{-}}\Big{(}|v-u|_{2,\tau}^{2}+h_{\tau}^{4}\|f-\Delta^{2}v\|_{0,\tau}^{2}\Big{)}$
$\lesssim|v-u|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{6,K^{{}^{\prime}}}^{2}.$
Step 3. Analogously, we need to construct an edge bubble function $b_{e}$ and an auxiliary polynomial $\phi\in H^{2}(T)$ such that
$\phi|_{e}=b_{e}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{|}_{e}\quad\text{and}\quad\phi|_{\overline{\partial T\setminus e}}=\partial_{\bm{n}}\phi|_{\overline{\partial T\setminus e}}=0,$
where $\overline{\partial T\setminus e}$ is the closure of $\partial T\setminus e$. Divide the triangle ${\tau}^{-}$ (resp. ${\tau}^{+}$) into two small triangles ${\tau}_{1}^{-},{\tau}_{2}^{-}$ (resp. ${\tau}_{1}^{+},{\tau}_{2}^{+}$) by connecting the midpoint of $e$ (denoted by $a_{e}$) and the vertex of ${\tau}^{-}$ (resp. ${\tau}^{+}$) opposite to the edge $e$. Assume the barycenter coordinate functions of the two small triangles at point $a_{e}$ are $b_{1}^{-}$, $b_{2}^{-}$ (resp. $b_{1}^{+}$, $b_{2}^{+}$). Let $b_{e}=(b_{1}^{-}b_{2}^{-}b_{1}^{+}b_{2}^{+})^{2}$ and $\phi=b_{e}E_{T}\Big{(}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{)}$. Then, we can check that these two functions satisfy the desired conditions. Furthermore, we have
$\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}^{2}\lesssim\Big{(}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]},\phi\Big{)}_{e},$
and by Green’s identity, there holds
$\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim\sum_{\tau=\tau^{-},\tau^{+}}\Big{(}(\Delta^{2}v,\,\phi)_{\tau}-a^{\tau}(v,\,\phi)+\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\,\partial_{\bm{n}^{\tau}}\phi\Big{)}_{0,e}\Big{)},$ (5.14)
where $\bm{n}^{\tau}$ is the unit outward normal vector of $\tau$ along $e$.
Since $\phi\in H_{0}^{2}(T)$, the equation (5.11) holds. Then, using (5.11), (5.14), the Cauchy-Schwarz inequality and the inverse inequalities of polynomials, we have
$\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim\sum\limits_{\tau=\tau^{+},\tau^{-}}\Big{(}(\Delta^{2}v-f,\,\phi)_{\tau}+a^{\tau}(u-v,\,\phi)\Big{)}+\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{0,e}$
$\lesssim\sum_{\tau=\tau^{+},\tau^{-}}\Big{(}\|f-\Delta^{2}v\|_{0,\tau}+h_{\tau}^{-2}|v-u|_{2,\tau}\Big{)}\|\phi\|_{0,\tau}+\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{0,e}\|\partial_{\bm{n}}\phi\|_{0,e}.$ (5.15)
Similar to (5.13), one has
$\|\phi\|_{0,\tau^{\pm}}\lesssim\Big{\|}E_{T}\Big{(}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{)}\Big{\|}_{0,\tau^{\pm}}\lesssim h_{K^{\pm}}^{1/2}\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}.$ (5.16)
Noting that $h_{K}\eqsim{h_{e}}$, we have by the trace inequality and the inverse inequalities of polynomials that
$\|\partial_{\bm{n}}\phi\|_{0,e}\lesssim h_{T}^{1/2}|\phi|_{2,T}+h_{T}^{-1/2}|\phi|_{1,T}\lesssim h_{e}^{-3/2}\|\phi\|_{0,T}.$ (5.17)
The desired result follows immediately from (5.6), (5.7), (5.15), (5.16) and (5.17). ∎
The following theorem shows the efficiency of the estimator globally, which is a direct consequence of the last theorem.
Theorem 5.3 (Lower bound).
Given $f\in H^{k-2}(\Omega)$ with $k\geq 2$, let $u$, $u_{h}$ and $\eta(u_{h})$ be given as in Theorem 5.1. Then there holds
$\eta(u_{h})\lesssim\|u-u_{h}\|_{h}+\mathcal{O}(h^{k-1}).$ (5.18)
Proof.
It follows from the definition of $\eta(u_{h})$ and Theorem 5.2 that
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})$
$+\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}+\sum_{K\in\mathcal{T}_{h}}\eta_{5,K}^{2}.$ (5.19)
From (3.13), we obtain
$\eta_{5,K}=h_{K}^{2}\|f-f_{h}\|_{0,K}\lesssim h^{k+2}\|f\|_{k-2}=\mathcal{O}(h^{k+2}).$ (5.20)
By the norm equivalence (3.17),
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$ $\lesssim|u-u_{h}|_{2,K}^{2}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$
$\lesssim|u-u_{h}|_{2,K}^{2}+S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})$
$\lesssim|u-u_{h}|_{2,K}^{2}+\eta_{4,K}^{2}.$ (5.21)
It follows from the norm equivalence in (3.17) and the triangle inequality that
$\eta_{4,K}$ $=h_{K}^{-1}\|\bm{\chi}(u_{h}-\Pi_{h}^{\Delta}u_{h})\|_{\ell^{2}}\eqsim|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}\lesssim|u-u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}.$
Here, by the triangle inequality, (3.8) and (2.4), we derive
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|\Pi_{h}^{\Delta}u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u|_{2,K}\lesssim|u-u_{h}|_{2,K}+h^{k-1}.$
Inserting (5.20) and (5.2) into (5.2), one has
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})+\mathcal{O}(h^{2k-2}).$ (5.22)
The proof is completed by the definition of the norm $\|\cdot\|_{h}$. ∎
5.3   Comparison with the analysis of Morley-type virtual elements
Ref. [13] provided the a posteriori error analysis for the biharmonic equation using the lowest-order Morley-type or $H^{2}$-nonconforming virtual elements. However, the local error estimators there do not account for the jumps of the second- and third-order derivatives across interior edges, specifically $\eta_{2,K}$ and $\eta_{3,K}$. As seen in the proof of Theorem 5.1, these additional terms arise from integrating by parts for $I_{2}$ in (5.4) and estimating the average term of $J_{3}$ in (5.5). For the Morley-type VEM, including $J_{i}$ ($i=1,2,3$) is unnecessary. Therefore, it is expected to avoid integration by parts of $I_{2}$ for the Morley-type VEM, which is achieved by means of a special property of the $H^{2}$-nonconforming interpolation operator described below.
Lemma 5.1.
Let $V_{k}^{2}(K)$ be the local $H^{2}$-nonconforming virtual element space [2, 30, 13]. Denote by $I_{K}^{2}$ the interpolation operator from $H^{2}(K)$ to $V_{k}^{2}(K)$. Then for any $v\in H^{2}(K)$ and $w\in V_{k}^{2}(K)$, there holds
$(\nabla^{2}(v-I_{K}^{2}v),\nabla^{2}w)_{K}=0,$
This implies that $I_{K}^{2}$ is also an $H^{2}$-elliptic projector from $H^{2}(K)$ to $V_{k}^{2}(K)$.
Proof.
See the proof of [21, Corollary 4.1]. ∎
Our analysis of the upper bound in Theorem 5.1 can be extended to the Morley-type VEM. Utilizing Lemma 5.1, we can avoid the tedious integration by parts for the term $I_{2}$ as follows:
$I_{2}=(f_{h},\phi-I_{h}^{2}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\Delta}u_{h},\phi-I_{h}^{2}\phi)=(f_{h},\phi-I_{h}^{2}\phi)=(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}^{2}\phi),$
where the last equality holds because $\Delta^{2}\Pi_{h}^{\Delta}u_{h}|_{K}\in\mathbb{P}_{k-4}(K)$ and $(I_{h}^{2}\phi,p)_{K}=(\phi,p)_{K}$ for $p\in\mathbb{P}_{k-4}(K)$. Consequently, there is no need to include the higher-order jump terms $\eta_{2,K}$ and $\eta_{3,K}$ in the global error estimator for the Morley-type VEMs. However, this does not apply to other types of VEMs for the biharmonic equation even for the $C^{1}$-continuous VEMs discussed in [16].
Another difference in the error estimators from [13] is the substitution of the local estimators $\eta_{0,e}(u_{h})$ and $\eta_{1,e}(u_{h})$ with $\eta_{0,e}(\Pi_{h}^{\Delta}u_{h})$ and $\eta_{1,e}(\Pi_{h}^{\Delta}u_{h})$. For the Morley-type VEM in [13], unlike the triangle inequality
$|u-u_{h}|_{2,K}\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-E_{h}u_{h}|_{2,K}$ (5.23)
used in (5.2), they utilize the following inequality
$|u-u_{h}|_{2,K}\leq|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-u_{h}|_{2,K}.$ (5.24)
The second term corresponds to the error estimator $\eta_{4,K}$. For the first term, since $\Pi_{h}^{\Delta}u_{h}|_{K}$ is a polynomial on $K$, they treat $\Pi_{h}^{\Delta}u_{h}$ as a finite element function on the virtual triangulation $\widehat{\mathcal{T}}$, defining $\widehat{E}_{h}:\mathbb{P}_{2}(\widehat{\mathcal{T}})\to H_{0}^{2}(\Omega)$ as the enriching operator for the FEM. They then decompose the first term as
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|u-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}.$ (5.25)
Let $\widehat{E}_{h}\Pi_{h}^{\Delta}:=\widetilde{E}_{h}$. The decomposition given by (5.24) and (5.25) can be written as
$|u-u_{h}|_{2,K}$ $\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}$
$\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}.$
The first term can be estimated as for the first term in (5.23), namely the first term in (5.2) for Theorem 5.1. By using the standard result for $\widehat{E}$, the third term is bounded by
$|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,h}^{2}=|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
For any $v\in V_{h}$, it is evident that $E_{h}\phi$ can be defined for $\phi=\Pi_{h}^{\Delta}v$. By the definition of $E_{h}$, we have $E_{h}v=E_{h}\Pi_{h}^{\Delta}v$, which along with Lemma 4.1 implies
$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
Therefore, we can perform the similar decomposition
$|u-u_{h}|_{2,K}$ $\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-E_{h}u_{h}|_{2,K}$
$\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-E_{h}u_{h}|_{2,K}.$
This approach allows us to derive the error estimators in [13] for the Morley VEMs of arbitrary order, not just the lowest order. However, for interior penalty methods, this may not be appropriate, as including the elliptic projector $\Pi_{h}^{\Delta}$ in $J_{1}$ (see (3.11)) is not feasible because $\|\cdot\|_{h}$ may not be a norm on $V_{h}$ in this context and $J_{1}(u-u_{h},u-u_{h})$ will introduce the a-priori information.
6 Numerical experiments
In this section, we present several benchmark tests for which the exact solution is known or at least its behavior is known in a qualitative sense. These tests illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM. Unless otherwise specified, we only consider the lowest-order element $k=2$. Since the VEM solution $u_{h}$ is not explicitly known inside the polygonal elements, we evaluate the projection error, which is quantified by
[…]
The first term in (3.9) will be disregarded when focusing solely on the $H^{2}$-semi-norm.
Standard adaptive algorithms based on the local mesh refinement can be written as loops of the form
${\bf SOLVE}\to{\bf ESTIMATE}\to{\bf MARK}\to{\bf REFINE}.$
Given an initial polygonal subdivision $\mathcal{T}_{0}$, to get $\mathcal{T}_{k+1}$ from $\mathcal{T}_{k}$ we first solve the VEM problem under consideration to get the numerical solution $u_{k}$ on $\mathcal{T}_{k}$. The error is then estimated by using $u_{k}$, $\mathcal{T}_{k}$ and the a posteriori error bound. The local error bound is used to mark a subset of elements in $\mathcal{T}_{k}$ for refinement. The marked polygons and possible more neighboring elements are refined in such a way that the subdivision meets certain conditions, for example, the resulting polygonal mesh is still shape regular.
In the implementation, all the initial meshes are generated by the PolyMesher package [23], which is a polygonal mesh generator based on the Centroidal Voronoi Tessellations. In the MARK step, We employ the Dörfler marking strategy with parameter $\theta\in(0,1)$ to select the subset of elements for refinement. In the REFINE step, it is usually time-consuming to write a mesh refinement function since we need to carefully design the rule for dividing the marked elements to get a refined mesh of high quality. We divide every polygonal element by connecting the midpoint of each edge to its barycenter, which requires that the barycenter is an internal point of each element. Moreover, for any two polygons $K_{1}$ and $K_{2}$ sharing a common edge $e$, we refine them together if $K_{1}$ is in the refinement set and one endpoint of $e$ is the hanging node of $K_{2}$, in order to avoid occurrence of short-edges. Such an implementation can be found in [26].
All examples are implemented in MATLAB R2019b. Our code is available from GitHub (https://github.com/Terenceyuyue/mVEM) as part of the mVEM package which contains efficient and easy-following codes for various VEMs published in the literature. The subroutine IPVEM_Zhao.m is used to compute the numerical solutions. The test script main_IPVEM_Zhao_avem.m verifies the convergence rates. The error estimator is computed by the subroutine IPVEM_Zhao_indicator.m.
6.1   Solution without large gradients
Example 6.1.
Let $\Omega=(0,1)^{2}$ and we select the load term $f$ such that the analytical solution of the problem (3.11) is
  • 1.
    homogeneous boundary value conditions:
    $u(x,y)=10x^{2}y^{2}(1-x)^{2}(1-y)^{2}\sin(\pi x).$
  • 2.
    or inhomogeneous boundary value conditions:
    $u(x,y)=10x^{2}y^{2}(1-x)^{2}(1-y)^{2}\sin(\pi x)+x^{2}+y^{2}.$
Fig. 1: The final adapted mesh and the exact and numerical solutions with homogeneous boundary values conditions
Fig. 2: (a) The estimator $\eta(u_{h},\mathcal{T}_{h})$ for adapted meshes; (b) The convergence orders
In this example, we employ the Dörfler marking strategy with a parameter of $\theta=0.4$ to select the subset of elements for refinement. The final adapted mesh and the solutions are displayed in Fig. 1 for the problem with homogeneous boundary values conditions. Since the exact solution is sufficiently smooth without large gradients, the initial mesh is refined almost everywhere. The estimator $\eta(u_{h},\mathcal{T}_{h})$ for all adapted meshes is plotted in Fig. 2a, showing a decrease throughout the adaptive procedure as expected. The convergence order is also displayed in Fig. 2b, where we observe the optimal first-order convergence rate for both the a priori error and the error estimator, consistent with theoretical predictions. Almost the same results are observed for this problem with inhomogeneous boundary values conditions. The final adapted mesh and the solutions are displayed in Fig. 3.
Fig. 3: The final adapted mesh and the exact and numerical solutions with inhomogeneous boundary values conditions
6.2   Solution with large gradients
Example 6.2.
Let $\Omega=(0,1)^{2}$ and we select the load term $f$ such that the analytical solution of the problem (3.11) is
$u(x,y)=xy(1-x)(1-y){\mathrm{exp}}\left(-1000((x-0.5)^{2}+(y-0.117)^{2})\right).$
Fig. 4: The initial and the final adapted meshes. (a) The initial mesh; (b) After 20 refinement steps; (c) After 30 refinement steps; (d) The zoomed mesh in (c)
Fig. 5: Convergence rates of the error $|u-\Pi_{h}^{\Delta}u_{h}|_{2,h}$ and the error estimator $\eta(u_{h})$
(a) Exact
(b) Numerical
Fig. 6: The exact and numerical solutions
For this example, we choose the parameter $\theta=0.6$. The initial mesh and the final adapted meshes after 20 and 30 refinement steps are presented in Fig. 4 (a-c), respectively. The detail of the last mesh is shown in Fig. 4 (d). Clearly, no small edges are observed. We plot the adaptive convergence order in Fig. 6, which confirms the theoretical prediction as in the last example. We also display the numerical and exact solutions in Fig. 5, from which we see the adaptive strategy correctly refines the mesh in a neighborhood of the singular point and there is a good level of agreement between the $H^{2}$ error and error estimator.
7 Conclusion
We developed a residual-type a posteriori error estimator for an interior penalty virtual element method applied to a Kirchhoff plate bending problem. A key step involves constructing an enriching operator to establish the reliability and efficiency of the a posteriori error bound. We presented an error estimate for the enriching operator that aligns with the interior penalty finite element method, as detailed in Lemma 4.1. Additionally, we provided an example estimate in Lemma 4.2 with a proof idea applicable to the medius error analysis.
CRediT authorship contribution statement
Fang Feng and Yue Yu collaborated closely to shape the conceptualization, methodology, and writing of this research. Additionally, Yue Yu took charge of implementing the discrete method in the study.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data sharing is not applicable to this article as no data was generated or analyzed.
Acknowledgements
Fang Feng was supported by the National Science Foundation for Young Scientists of China (No. 12401528) and also partially supported by Central Government Special Fund for Basic Scientific Research Business Expenses of Colleges and Universities, No. 30924010837. Yue Yu was partially supported by the National Science Foundation for Young Scientists of China (No. 12301561).

### v4

@@
 Abstract
-A residual-type a posteriori error estimation is developed for an interior penalty virtual element method (IPVEM) to solve a Kirchhoff plate bending problem with inhomogeneous boundary value conditions.
-The computable error estimator is incorporated.
-We derive the reliability and efficiency of the a posteriori error bound by constructing an enriching operator and establishing some related error estimates that align with interior penalty finite element methods.
+In this paper, we develop a residual-type a posteriori error estimation for an interior penalty virtual element method (IPVEM) for the Kirchhoff plate bending problem.
+Building on the work in [18], we adopt a modified discrete variational formulation that incorporates the $H^{1}$-elliptic projector in the jump and average terms.
+This allows us to simplify the numerical implementation by including the $H^{1}$-elliptic projector in the computable error estimators.
+We derive the reliability and efficiency of the a posteriori error bound by constructing an enriching operator and establishing some related error estimates that align with $C^{0}$-continuous interior penalty finite element methods.
+As observed in the a priori analysis, the interior penalty virtual elements exhibit similar behaviors to $C^{0}$-continuous elements despite its discontinuity.
+This observation extends to the a posteriori estimate since we do not need to account for the jumps of the function itself in the discrete scheme and the error estimators.
 As an outcome of the error estimator, an adaptive VEM is introduced by means of the mesh refinement strategy with the one-hanging-node rule.
-Numerical results on several benchmark tests confirm the robustness of the proposed error estimator and show the efficiency of the resulting adaptive VEM.
+Numerical results from several benchmark tests confirm the robustness of the proposed error estimators and show the efficiency of the resulting adaptive VEM.
 keywords:
@@
 Recently, the interior penalty virtual element method (IPVEM) has been explored in [30] for the biharmonic equation.
-Compared with the conforming or nonconforming VEMs for the biharmonic problem, the IPVEM has advantages in reducing the number of degrees of freedom (DoFs), which can be regarded as a combination of the virtual element space and discontinuous Galerkin scheme, since the resulting global discrete space is not $C^{0}$-continuous and an interior penalty formulation is adopted to enforce the continuity of the solution.
+Compared with the conforming or nonconforming VEMs for the biharmonic problem, the IPVEM has advantages in reducing the number of degrees of freedom (DoFs), which can be regarded as a combination of the virtual element space and discontinuous Galerkin scheme.
 This method was adapted for the fourth-order singular perturbation problem in [27], with some adaptations to the original space.
@@
 In contrast, Ref.
-[18] employs the original IPVEM formulation from [29] to address the fourth-order singular perturbation problem (2.1), incorporating techniques from the modified Morley finite element method as described in [25].
-On the other hand, due to the large flexibility of the meshes in VEMs, the a posteriori error analysis of VEMs is well worth studying.
+[18] employs the original IPVEM formulation from [30] to address the fourth-order singular perturbation problem, incorporating techniques from the modified Morley finite element method as described in [25].
+Another point to be mentioned is that due to the large flexibility of the meshes in VEMs, there is no exceptional treatment for hanging nodes, which has great superiority in mesh refinement.
+Hence, a posteriori error analysis of VEMs is well worth studying.
 For the fourth-order problem, a residual-based a posteriori error estimate for the $C^{1}$-continuous VEM was developed in [16], addressing both the reliability and efficiency of the error bound.
@@
 This reduction is facilitated by the special property of the interpolation operator for the Morley virtual element (see Lemma 5.1).
-In this paper, we consider the a posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates.
-We provide a detailed description for implementing the IPVEM with inhomogeneous boundary conditions.
-For the sake of the construction of error estimators, unlike the previous interior penalty formulations in [29, 18], we substitute the $H^{1}$-elliptic projection operator in the jump and average terms for second-order derivatives with an $H^{2}$-elliptic projection operator.
-Following the idea in [21], we construct an enriching operator to establish the reliability and efficiency of the a posteriori error bound.
-This enriching operator inherits the beneficial properties of the operators described in [13] for the lowest-order Morley virtual element and is applicable to arbitrary-order cases.
+Recently, the medius error estimates for the interior penalty virtual element method for the biharmonic equation are proposed in [29].
+In this paper, we consider a posteriori error estimation for the interior penalty virtual element method applied to Kirchhoff plate bending problems.
+Unlike the jump and average terms in [30], we include the $H^{1}$-elliptic projector $\Pi_{h}^{\nabla}$ in the penalty terms, as proposed in [18].
+This modification enables the design of computable error estimators that incorporate the $H^{1}$-elliptic projector, thereby simplifying the implementation process.
+Following the approach in [21], we construct an enriching operator to establish the reliability of the a posteriori error estimators.
+We then develop a residual-type a posteriori error estimation for the modified IPVEM applied to Kirchhoff plates.
+It is worth noting that although the interior penalty virtual element method is discontinuous, as observed in the a priori error estimate, the interior penalty virtual elements exhibit behavior similar to $C^{0}$-continuous elements.
+This “quasi-$C^{0}$ continuity” can be extended to the a posteriori estimate since we do not need to account for the jumps of the function itself in the discrete scheme or the error estimators.
 The remainder of the paper is structured as follows.
 We begin by introducing the continuous variational problem and presenting some useful results in VEM analysis in Section 2.
-Section 3 details the interior penalty virtual element method, including the virtual element space, degrees of freedom, and the discrete bilinear form, with a particular focus on addressing inhomogeneous boundary value conditions not covered in previous references.
+Section 3 describes the interior penalty virtual element method.
+We also incorporate inhomogeneous boundary conditions not covered in previous references in the section for numerical experiments.
+In order to simplify the implementation, we add the elliptic projector $\Pi_{h}^{\nabla}$ for all $v,w$ in the jump and average terms.
 In Section 4, we construct an enriching operator and present the corresponding error estimates.
 Section 5 develops the a posteriori analysis for the IPVEM, providing both upper and lower bounds.
+We do not include the jumps of the function itself in the computable posteriori error estimators despite the discontinuity of the interior penalty virtual element.
 Meanwhile, we make a comparison with the analysis of Morley-type virtual elements in [13].
-An adaptive IPVEM is presented in Section 6 with several benchmark tests performed to illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM.
+An adaptive IPVEM is presented in Section 6 with several benchmark tests performed to illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive IPVEM.
 2 The continuous variational problem
 Let $\Omega$ be a bounded polygonal domain of $\mathbb{R}^{2}$ with boundary $\partial\Omega$. For $f\in L^{2}(\Omega)$, we consider the following boundary value problem of the Kirchhoff plate:
-$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=g_{D},\quad\dfrac{\partial u}{\partial\bm{n}}=g_{N}&{\rm on}~{}~{}\partial\Omega.\end{array}\right.$ (2.1)
-where $\bm{n}=(n_{1},n_{2})$ is the unit outer normal to $\partial\Omega$, $\Delta$ is the standard Laplacian operator, and $g_{D}\in H^{3/2}(\partial\Omega)$ and $g_{N}\in H^{1/2}(\partial\Omega)$ are the Dirichlet data.
-Throughout the article we assume $u\in H^{k+1}(\Omega)$ with $k\geq 2$
+$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=0,\quad\dfrac{\partial u}{\partial\bm{n}}=0&{\rm on}~{}~{}\partial\Omega,\end{array}\right.$ (2.1)
+where $\bm{n}=(n_{1},n_{2})^{T}$ is the unit outer normal to $\partial\Omega$, $\Delta$ is the standard Laplacian operator. Throughout the article we assume $u\in H^{k+1}(\Omega)$ with $k\geq 2$.
 […]
@@
 From the definition of the averages and jumps, by a direct manipulation, one has $[vw]=[v]\{w\}+\{v\}[w]$ for $e\in\mathcal{E}_{h}^{0}$ and $[vw]=vw=\{v\}[w]=[v]\{w\}$ for a boundary edge, leading to
-$\sum_{e\in\mathcal{E}_{h}}\int_{e}[vw]\mathrm{d}s=\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}[v]\{w\}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\int_{e}\{v\}[w]\mathrm{d}s,$ (2.2)
-which will be used frequently later.
+$\sum_{e\in\mathcal{E}_{h}}\int_{e}[vw]\mathrm{d}s=\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}[v]\{w\}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\int_{e}\{v\}[w]\mathrm{d}s.$ (2.2)
 Moreover, for any two quantities $a$ and $b$, “$a\lesssim b$” indicates “$a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and “$a\eqsim b$” abbreviates “$a\lesssim b\lesssim a$”.
@@
 $a(u,v)=(f,v),\quad v\in V,$ (2.3)
-where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$.
-To keep the flexibility of meshes, we will work on meshes satisfying the following assumption (cf. [10, 14]):
+where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$. To keep the flexibility of meshes, we will work on meshes satisfying the following assumption (cf. [10, 14]):
   1. H.
@@
 As shown in [14], this condition covers the usual geometric assumptions frequently used in the context of VEMs.
-According to the standard Dupont-Scott theory (cf.
-[7]), for all $v\in H^{l}(K)$ ($0\leq l\leq k$) there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
+Under this geometric assumption, we can establish some fundamental results in VEM analysis as used in [21], with the constants hidden in the symbol $\lesssim$ depending only on the shape regularity and quasi-uniformity of the virtual triangulation ${\mathcal{T}_{K}}$.
+According to the standard Dupont-Scott theory (cf. [7]), for all $v\in H^{l}(K)$ ($0\leq l\leq k$) there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
 $|v-q|_{m,K}\lesssim h_{K}^{l-m}|v|_{l,K},\quad m\leq l.$ (2.4)
-The following trace inequality [7] and the Sobolev inequality [6] are very useful for our forthcoming analysis.
-Lemma 2.1 (Trace inequality).
+The following inequalities are very useful for our forthcoming analysis.
+Lemma 2.1.
 For any $K\in\mathcal{T}_{h}$, it holds
-$\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K).$ (2.5)
-Lemma 2.2 (Sobolev inequality).
-For any $K\in\mathcal{T}_{h}$, it holds
-$\|v\|_{\infty,K}\lesssim h_{K}|v|_{2,K}+|v|_{1,K}+h_{K}^{-1}\|v\|_{0,K},\quad v\in H^{2}(K).$ (2.6)
+  • 1.
+    Trace inequality [7]:
+    $\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K),$ (2.5)
+  • 2.
+    Sobolev inequality [6]:
+    $\|v\|_{\infty,K}\lesssim h_{K}|v|_{2,K}+|v|_{1,K}+h_{K}^{-1}\|v\|_{0,K},\quad v\in H^{2}(K),$ (2.6)
+  • 3.
+    Poincaré-Friedriches inequality [5]:
+    $\|v\|_{0,K}\lesssim h_{K}|v|_{1,K}+\Big{|}\int_{\partial K}v\,{\rm d}s\Big{|},\quad v\in H^{1}(K).$ (2.7)
 3 The interior penalty virtual element method
-This section reviews the interior penalty virtual element space developed in [29] and details the treatment of inhomogeneous boundary value conditions for the IPVEM.
+This section reviews the interior penalty virtual element space developed in [30].
 3.1   The IP virtual element spaces
 In the construction, the authors in [30] first introduced a $C^{1}$-conforming virtual element space
-$\widetilde{V}_{k+2}(K)=\left\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K\right\},\quad k\geq 2.$
+$\widetilde{V}_{k+2}(K)=\Big{\{}v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),e\subset\partial K\Big{\}},\quad k\geq 2.$
 This local space can be equipped with the following degrees of freedom (DoFs) (cf. [11, 17]):
@@
     $\tilde{\chi}_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k}(K).$ (3.5)
+Referring to [30, Lemma 3.4], we have the following lemma.
+Lemma 3.1 (Inverse inequality).
+For any $K\in\mathcal{T}_{h}$, for $m=1,2$ and $0\leq s\leq m$, there holds
+$|v_{h}|_{m,K}\lesssim h_{K}^{s-m}|v_{h}|_{s,K},\quad v_{h}\in\widetilde{V}_{k+2}(K).$ (3.6)
 Given $v_{h}\in\widetilde{V}_{k+2}(K)$, the usual definition of the $H^{1}$-elliptic projection $\Pi_{K}^{\nabla}v_{h}\in$ $\mathbb{P}_{k}(K)$ is described by the following equations:
 $(\nabla\Pi_{K}^{\nabla}v_{h},\nabla q)_{K}&=(\nabla v_{h},\nabla q)_{K},\quad q\in\mathbb{P}_{k}(K),\\ \sum\limits_{z\in\mathcal{V}_{K}}\Pi_{K}^{\nabla}v_{h}(z)&=\sum\limits_{z\in\mathcal{V}_{K}}v_{h}(z),\end{aligned}\right.$ (3.7)
-where $\mathcal{V}_{K}$ is the set of the vertices of $K$. By checking the right-hand side of the integration by parts formula:
-$(\nabla v_{h},\nabla q)_{K}=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}\int_{e}v_{h}\partial_{\bm{n}_{e}}q\,\mathrm{d}s.$
-However, the goal of the IPVEM is to make $\Pi_{K}^{\nabla}v_{h}$ computable by only using the DoFs of $H^{1}$-conforming virtual element spaces given by (cf. [3, 1])
+where $\mathcal{V}_{K}$ is the set of the vertices of $K$. By checking the right-hand side of the integration by parts formula
+$(\nabla v_{h},\nabla q)_{K}=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}\int_{e}v_{h}\partial_{\bm{n}_{e}}q\,\mathrm{d}s,$
+one can find that this elliptic projection can be computed using the DoFs in (3.1)–(3.5).
+However, the goal of the IPVEM is to make $\Pi_{K}^{\nabla}v_{h}$ computable by only using the DoFs of $H^{1}$-conforming virtual element spaces given by (cf.
+[3, 1])
 $V_{h}^{1,c}(K):=\{v\in H^{1}(K):\Delta v|_{K}\in\mathbb{P}_{k-2}(K)~{}~{}{\text{in}}~{}~{}K,\quad v|_{\partial K}\in\mathbb{B}_{k}(\partial K)\},$
@@
 $V_{h}^{1,c}=\{v\in C(\overline{\Omega}):v|_{K}\in V_{h}^{1,c}(K),\,K\in\mathcal{T}_{h}\}\cap H_{0}^{1}(\Omega).$
-To do so, Ref.
-[29] considered the approximation of the right-hand side by some numerical formula.
-In view of the $\mathbb{P}_{k}$ accuracy, namely $\Pi_{K}^{\nabla}v_{h}=v_{h}$ for $v_{h}\in\mathbb{P}_{k}(K)$, the modified $H^{1}$-projection is defined as
+To do so, Ref. [30] considered an approximation of the right-hand side by some numerical formula. The modified $H^{1}$-projection is defined as
 $(\nabla\Pi_{K}^{\nabla}v_{h},\nabla q)_{K}&=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}Q_{2k-1}^{e}(v_{h}\partial_{\bm{n}_{e}}q),\quad q\in\mathbb{P}_{k}(K),\\ \sum_{z\in\mathcal{V}_{K}}\Pi_{K}^{\nabla}v_{h}(z)&=\sum_{z\in\mathcal{V}_{K}}v_{h}(z),\end{aligned}\right.$
@@
 where $(\omega_{i},\bm{x}_{i}^{e})$ are the $(k+1)$ Gauss-Lobatto quadrature weights and points with $\bm{x}_{0}^{e}$ and $\bm{x}_{k}^{e}$ being the endpoints of $e$.
-With the help of $\Pi_{K}^{\nabla}$, we can employ the standard enhancement technique by substituting the redundant moments of $v$ with those of $\Pi_{K}^{\nabla}v$ [1].
+With the help of $\Pi_{K}^{\nabla}$, we can employ the standard enhancement technique by substituting the redundant DoFs of $v$ with those of $\Pi_{K}^{\nabla}v$ [1, 30].
 The local interior penalty space is defined as
-$V_{k}(K)=\{v\in W_{k}(K):\tilde{\chi}^{g}(v)=\tilde{\chi}^{g}(\Pi_{K}^{\nabla}v),\tilde{\chi}^{n}(v)=\tilde{\chi}^{n}(\Pi_{K}^{\nabla}v)\},$
-where
-$W_{k}(K):=\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k-2}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~{}~{}e\subset\partial K,$
-$\hskip 113.81102pt(v,q)_{K}=(\Pi_{K}^{\nabla}v,q)_{K},\quad q\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K)\}.$
+$V_{k}(K)=\{v\in\widetilde{V}_{k+2}(K):\tilde{\chi}^{g}(v)=\tilde{\chi}^{g}(\Pi_{K}^{\nabla}v),\tilde{\chi}^{n}(v)=\tilde{\chi}^{n}(\Pi_{K}^{\nabla}v),$
+$\hskip 113.81102pt(v,q)_{K}=(\Pi_{K}^{\nabla}v,q)_{K}\quad q\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K)\}\quad k\geq 2,$
 The associated DoFs are given by
@@
 […]
+Denote by $I_{h}:H^{2}(\Omega)\to V_{h}$ the interpolation operator. As shown in Lemma 3.11 of [30], we can derive the following interpolation error estimate.
+Lemma 3.2.
+For any $v\in H^{\ell}(K)$ with $12\leq\ell\leq k+$ 1, it holds
+$|v-I_{h}v|_{m,K}\lesssim h_{K}^{\ell-m}|v|_{\ell,K},\quad K\in\mathcal{T}_{h},\quad m=0,1,2.$ (3.8)
 As usual, we can define the $H^{2}$-projection operator $\Pi_{K}^{\Delta}:V_{k}(K)\to\mathbb{P}_{k}(K)$ by finding the solution $\Pi_{K}^{\Delta}v\in\mathbb{P}_{k}(K)$ of
@@
 $\widehat{v}=\frac{1}{|\partial K|}\int_{\partial K}v\mathrm{d}s.$
-Choosing $q=\Pi_{K}^{\Delta}v$ in (3.7), we can easily observe
-$|\Pi_{K}^{\Delta}v|_{2,K}\lesssim|v|_{2,K},\qquad v\in H^{2}(K).$ (3.8)
-3.2   The IPVEM with inhomogeneous boundary value conditions
-Let $v\in H^{2}(K)$. The integration by parts gives
-$(f,v)_{K}=(\Delta^{2}u,v)_{K}=(\nabla^{2}u,\nabla^{2}v)_{K}+\int_{\partial K}\Big{(}\frac{\partial\Delta u}{\partial\bm{n}_{K}}v+\frac{\partial^{2}u}{\partial\bm{n}_{K}\partial\bm{t}_{K}}\frac{\partial v}{\partial\bm{t}_{K}}\Big{)}\mathrm{d}s-\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v}{\partial\bm{n}_{K}}\mathrm{d}s.$
-If $v$ is $C^{0}$-continuous, the second term on the right-hand side naturally vanishes when we sum the equation over all polygons in $\mathcal{T}_{h}$.
-However, when $v=v_{h}$ is a function in the interior penalty virtual element space $V_{h}$, which is not continuous, the second term contributes to the summation.
-The IPVEM in [29] neglects this contribution as it introduces a consistency error with a desired convergence order as shown in [29, 18], which gives
-$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s.$
-Using the formula in (2.2) and noting that $u\in H^{3}(\Omega)$, we get
-$\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s$ $=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}\Big{[}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\{}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{\}}\mathrm{d}s$
-$=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
-yielding
-$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
-To ensure both the symmetry and $C^{1}$-continuity of the discrete variational problem, we rewrite it as
-$(f,v_{h})-\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{\partial}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
-$\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
-$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
-where we have used the fact that $[u]_{e}=0$ and $[\frac{\partial u}{\partial\bm{n}_{e}}]_{e}=0$ for $e\in\mathcal{E}_{h}^{0}$.
-Following the standard procedure in the VEMs, we define the discrete bilinear form as
-$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$
+Lemma 3.3.
+For every $v\in V_{k}(K)$, there hold
+$|\Pi_{K}^{\nabla}v|_{m,K}\lesssim|v|_{m,K}\quad\mbox{and}\quad|\Pi_{K}^{\Delta}v|_{m,K}\lesssim|v|_{m,K},$
+where $m=0,1,2$.
+Proof.
+The first inequality is presented in Corollary 3.7 of [30].
+For the second one, it is simple to find that $|\Pi_{K}^{\Delta}v|_{2,K}\lesssim|v|_{2,K}$.
+For $m=1$, by the triangle inequality, the Poincaré-Friedrichs inequality (2.7), the boundedness of $|\Pi_{K}^{\Delta}v|_{2,K}$ and the inverse inequality (3.6), we derive
+$|\Pi_{K}^{\Delta}v|_{1,K}$ $\leq|v-\Pi_{K}^{\Delta}v|_{1,K}+|v|_{1,K}\lesssim h_{K}|v-\Pi_{K}^{\Delta}v|_{2,K}+|v|_{1,K}$
+$\lesssim h_{K}|v|_{2,K}+|v|_{1,K}\lesssim|v|_{1,K}.$
+The case of $m=0$ can be deduced in the similar manner. ∎
+3.2   The discrete IP bilinear form
+Following the standard procedure in the VEMs (namely, we need to consider the computability, the $k$-consistency (3.17) and the stability (3.18)), we define the discrete bilinear form as
+$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+h_{K}^{-2}S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$
 with
-$S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w)=h_{K}^{-2}\sum\limits_{i=1}^{n_{K}}\chi_{i}(v-\Pi_{K}^{\Delta}v)\chi_{i}(w-\Pi_{K}^{\Delta}w),$
-where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. Let
-$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[v][w]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$ (3.9)
-with $\lambda_{e}\geq 1$ being some edge-dependent parameter, and the additional two terms are given by
-$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
-$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
-Here and bellows, we define the piecewise $H^{2}$ projector $\Pi_{h}^{\Delta}$ by setting $\Pi_{h}^{\Delta}|_{K}=\Pi_{K}^{\Delta}$ for all $K\in\mathcal{T}_{h}$.
-The IPVEM for (2.1) with inhomogeneous boundary value conditioins can be described as follows: Find $u_{h}\in V_{h}^{g}=\{v\in V_{h}:v=I_{h}g_{D}~{}{\rm on}~{}\partial\Omega\}$ such that
+$S^{K}(v,w)=\sum\limits_{i=1}^{n_{K}}\chi_{i}(v)\chi_{i}(w),$
+where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. As was done in [18], we define
+$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$ (3.10)
+$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
+$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
+Here and below, we define the piecewise $H^{1}$ and $H^{2}$ projectors $\Pi_{h}^{\nabla}$ and $\Pi_{h}^{\Delta}$ by setting $\Pi_{h}^{\nabla}|_{K}=\Pi_{K}^{\nabla}$ and $\Pi_{h}^{\Delta}|_{K}=\Pi_{K}^{\Delta}$ for all $K\in\mathcal{T}_{h}$.
+The bilinear form is
+$\mathcal{A}_{h}(u_{h},v_{h})=a_{h}(u_{h},v_{h})+J_{1}(u_{h},v_{h})+J_{2}(u_{h},v_{h})+J_{3}(u_{h},v_{h}),$ (3.11)
+where $a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w)$. The discrete scheme of IPVEM solving problem (2.1) is to find $u_{h}\in V_{h}$ such that
 $\mathcal{A}_{h}(u_{h},v_{h})=F_{h}(v_{h}),\quad v_{h}\in V_{h},$ (3.12)
-where $g_{D}$ has been extended to the whole domain. The bilinear form is
-$\mathcal{A}_{h}(u_{h},v_{h})=a_{h}(u_{h},v_{h})+J_{1}(u_{h},v_{h})+J_{2}(u_{h},v_{h})+J_{3}(u_{h},v_{h}),$ (3.11)
-where $a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w)$. The right-hand side is
-$F_{h}(v_{h})=\langle f_{h},v_{h}\rangle+\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{(}-\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}+\frac{\lambda_{e}^{3}}{|e|^{3}}[u][v_{h}]+\frac{\lambda_{e}}{|e|}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{)}\mathrm{d}s,$
+The right-hand side is
+$F_{h}(v_{h}):=\langle f_{h},v_{h}\rangle,$
 where
 $\langle f_{h},v_{h}\rangle:=\sum\limits_{K\in\mathcal{T}_{h}}\int_{K}f\,\Pi_{0,K}^{k}v_{h}\mathrm{d}x=\sum\limits_{K\in\mathcal{T}_{h}}\int_{K}\Pi_{0,K}^{k}f\,v_{h}\mathrm{d}x,$ (3.13)
-with $\Pi_{0,K}^{k}$ being the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. Then for $f\in H^{k-2}(\Omega)$ with $k\geq 2$, we have
+with $\Pi_{0,K}^{k}$ being the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. For $f\in H^{k-2}(\Omega)$ with $k\geq 2$, we have
 $\langle f-f_{h},v_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}(f,v_{h}-\Pi_{0,K}^{k}v_{h})=\sum_{K\in\mathcal{T}_{h}}(f-\Pi_{0,K}^{k}f,v_{h}-\Pi_{0,K}^{k}v_{h})$
@@
 $\|f-f_{h}\|_{V^{\prime}_{h}}\lesssim h^{k}\|f\|_{k-2}.$ (3.14)
+Remark 3.1.
+As observed in [18], despite the discontinuity, the interior penalty virtual elements display behaviors similar to $C^{0}$-continuous elements in the a priori estimate.
+We will demonstrate that this observation also holds for the a posteriori estimate.
+Therefore, we do not include the penalty term
+$\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[v][w]\mathrm{d}s$
+in the discrete scheme and introduce the error estimator
+$\eta_{0,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{0,e}^{2}\Big{)}^{1/2},\qquad\eta_{0,e}=\frac{1}{|e|^{3/2}}\|[u_{h}]\|_{0,e}$ (3.15)
+in the a posteriori estimate, as will be discussed later.
 In what follows, we define
 $\|w\|_{h}^{2}:=|w|_{2,h}^{2}+J_{1}(w,w).$ (3.16)
-With the mesh-dependent parameter $\lambda_{e}$ chosen as in [18], we can deduce the stability result by using the same arguments. We omit the details with the results described as follows.
+Referring to [18], the mesh-dependent parameter $\lambda_{e}\eqsim 1$. Still, we can deduce the stability result by using the same arguments. We omit the details with the results described as follows.
     $k$-consistency: for all $v\in V_{k}(K)$ and $q\in\mathbb{P}_{k}(K)$, it holds that
@@
     for all $v\in V_{k}(K)$.
-For later use, we provide the following norm equivalence with the proof given in [29].
+For later use, we provide the following norm equivalence.
 Lemma 3.4.
-For $m=0,1,2$, there holds
-$|v-\Pi_{K}^{\Delta}v|_{m,K}\eqsim h_{K}^{1-m}\|\bm{\chi}(v-\Pi_{K}^{\Delta}v)\|_{\ell^{2}},\qquad v\in V_{k}(K).$ (3.17)
+For every $v\in V_{k}(K)$, there hold
+$|v-\Pi_{K}^{\Delta}v|_{2,K}\eqsim h_{K}^{-1}\|\bm{\chi}(v-\Pi_{K}^{\Delta}v)\|_{\ell^{2}},$ (3.19)
+$|v-\Pi_{K}^{\nabla}v|_{1,K}\eqsim\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}},$ (3.20)
+$h_{K}^{-1}\|v-\Pi_{K}^{\nabla}v\|_{0,K}\eqsim\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}}.$ (3.21)
+Proof.
+The first and second equations are presented in [27, Lemma 4.1].
+Following the approach in [21], it is straightforward to derive the lower bound estimate $\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}}\lesssim h_{K}^{-1}\|v-\Pi_{K}^{\nabla}v\|_{0,K}$ through direct calculations.
+For the upper bound estimate, according to the boundedness of $\Pi_{h}^{\nabla}$ and $\Pi_{h}^{\Delta}$ in Lemma 3.3 and the Poincaré-Friedrichs inequality (2.7), one has
+$\|v-\Pi_{h}^{\nabla}v\|_{0,K}$ $\leq\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\Delta}v-\Pi_{h}^{\nabla}v\|_{0,K}$
+$=\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\nabla}(\Pi_{h}^{\Delta}v-v)\|_{0,K}\lesssim\|v-\Pi_{h}^{\Delta}v\|_{0,K}$
+$\lesssim h_{K}|v-\Pi_{h}^{\Delta}v|_{1,K}\leq h_{K}|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}|\Pi_{h}^{\Delta}(\Pi_{h}^{\nabla}v-v)|_{1,K}$
+$\lesssim h_{K}|v-\Pi_{h}^{\nabla}v|_{1,K}.$ (3.22)
+This along with the second formula yields the desired result. ∎
 4 Enriching operator
 This section focuses on proposing and analyzing an enriching operator $E_{h}$ from $V_{h}$ to $H^{2}$-conforming virtual element spaces.
-The enriching operator is a special kind of quasi-interpolation operator, which connects a finite dimensional space to another one with higher order regularity.
+The enriching operator is a special kind of quasi-interpolation operator, which connects a finite-dimensional space to another one with higher-order regularity.
 We refer to [5, 9, 19, 22, 20, 8] and the references therein for the construction and application of such operators in finite element methods.
@@
     $\chi_{K}(v)=|K|^{-1}(m_{K},v)_{K},\quad m_{K}\in\mathbb{M}_{k-4}(K).$
-Note that the above moments vanish in the lowest order case $k=2$.
+Note that the above moments vanish in the lowest-order case $k=2$.
 The global virtual element spaces are then defined as
 $V_{h}^{2,c}:=\Big{\{}v\in H_{0}^{2}(\Omega)\cap C^{1}(\overline{\Omega}):v|_{K}\in V_{k}^{2,c}(K),~{}~{}K\in\mathcal{T}_{h}\Big{\}}.$
-Let ${\mathcal{E}_{a}^{0}}$ and $\mathcal{E}_{e}^{0}$ denote the interior vertices and edges of $\mathcal{T}_{h}$, respectively. The global dual space is defined as
+The global dual space is defined as
 $\mathcal{N}=\text{span}\Big{\{}\mathcal{N}_{a},\mathcal{N}_{\nabla},\mathcal{N}_{e}^{k-4},\mathcal{N}_{n_{e}}^{k-3},\mathcal{N}_{K}^{k-4}\Big{\}}$
@@
     $\mathcal{N}_{a}$: the values at the interior vertices of $\mathcal{T}_{h}$,
-    $\mathcal{N}_{a}^{z}(v)=v(z),\quad z\in\mathcal{E}_{a}^{0}.$
+    $\mathcal{N}_{a}^{z}(v)=v(z),\quad z\in\mathcal{V}_{h}^{0}.$
   • 2.
     $\mathcal{N}_{\nabla}$: the gradient values at interior vertices of $\mathcal{T}_{h}$,
-    $\mathcal{N}_{\nabla}^{z}(v)=h_{z}\nabla v(z),\quad z\in\mathcal{E}_{a}^{0}.$
+    $\mathcal{N}_{\nabla}^{z}(v)=h_{z}\nabla v(z),\quad z\in\mathcal{V}_{h}^{0}.$
   • 3.
     ${\mathcal{N}_{e}^{k-4}}$: the moments of $v$ on interior edges up to degree $k-4$,
-    $\mathcal{N}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e),\quad e\in\mathcal{E}_{e}^{0}.$
+    $\mathcal{N}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e),\quad e\in\mathcal{E}_{h}^{0}.$
   • 4.
     $\mathcal{N}_{n_{e}}^{k-3}$: the moments of $\partial_{\bm{n}}v$ on interior edges up to degree $k-3$,
-    $\mathcal{N}_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e),\quad e\in\mathcal{E}_{e}^{0}.$
+    $\mathcal{N}_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e),\quad e\in\mathcal{E}_{h}^{0}.$
   • 5.
@@
     For the values at interior vertices,
-    $\mathcal{N}_{a}^{z}(E_{h}\phi)=(E_{h}\phi)(z):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
+    $\mathcal{N}_{a}^{z}(E_{h}\phi)=(E_{h}\phi)(z):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}(z)\quad z\in\mathcal{V}_{h}^{0}.$
   2. 2.
     For the gradient values at interior vertices,
-    $\mathcal{N}_{\nabla}^{z}(E_{h}\phi):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}h_{z}\nabla\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
+    $\mathcal{N}_{\nabla}^{z}(E_{h}\phi):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}h_{z}\nabla\Pi_{h}^{\nabla}\phi|_{K^{\prime}}(z)\quad z\in\mathcal{V}_{h}^{0}.$
   3. 3.
     For the moments of $v$ on interior edges,
-    $\mathcal{N}_{e}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}|e|^{-1}\int_{e}m_{e}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}\mathrm{d}s,\quad m_{e}\in\mathbb{M}_{k-4}(e).$
+    $\mathcal{N}_{e}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}|e|^{-1}\int_{e}m_{e}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}\mathrm{d}s,\quad m_{e}\in\mathbb{M}_{k-4}(e).$
   4. 4.
     For the moments of ${\partial_{\bm{n}}}v$ on interior edges,
-    $\mathcal{N}_{n_{e}}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}\int_{e}m_{e}\partial_{\bm{n}}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}{\text{d}}s,\quad m_{e}\in\mathbb{M}_{k-3}(e).$
+    $\mathcal{N}_{n_{e}}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}\int_{e}m_{e}\partial_{\bm{n}}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}{\text{d}}s,\quad m_{e}\in\mathbb{M}_{k-3}(e).$
   5. 5.
     For the moments on element $K$,
-    $\mathcal{N}_{K}(E_{h}\phi):=|K|^{-1}\int_{K}m_{K}\Pi_{h}^{\Delta}\phi|_{K}{\mathrm{d}}x,\quad m_{K}\in\mathbb{M}_{k-4}(K).$
+    $\mathcal{N}_{K}(E_{h}\phi):=|K|^{-1}\int_{K}m_{K}\Pi_{h}^{\nabla}\phi|_{K}{\mathrm{d}}x,\quad m_{K}\in\mathbb{M}_{k-4}(K).$
 Lemma 4.1.
 For any $v\in V_{h}$ and $E_{h}v$ defined above, there holds
-$|v-E_{h}v|_{2,h}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)},$ (4.2)
+$|v-E_{h}v|_{2,h}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}.$ (4.2)
+In addition,
+$\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v),$ (4.3)
+$\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}.$ (4.4)
 Proof.
-According to the triangle inequality
-$|v-E_{h}v|_{2,h}\leq|v-\Pi_{h}^{\Delta}v|_{2,h}+|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}$
-and the norm equivalence (3.17), it reduces to bound the second term on the right-hand side.
-Since $\Pi_{h}^{\Delta}v-E_{h}v\in V_{k}^{2,c}(K)$, by the inverse inequality and norm equivalence in [21], one gets
-$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,K}\lesssim h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\lesssim h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}},\qquad K\in\mathcal{T}_{h}.$
+Step 1: We first prove (4.2). Consider the triangle inequality
+$|v-E_{h}v|_{2,h}\leq|v-\Pi_{h}^{\nabla}v|_{2,h}+|\Pi_{h}^{\nabla}v-E_{h}v|_{2,h}.$ (4.5)
+Using the inverse inequality (3.6) and the norm equivalence (3.20), we obtain
+$|v-\Pi_{h}^{\nabla}v|_{2,h}$ $=\sum_{K\in\mathcal{T}_{h}}|v-\Pi_{h}^{\nabla}v|_{2,K}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}|v-\Pi_{h}^{\nabla}v|_{1,K}$
+$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}(S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v))^{1/2}.$
+Thus, it reduces to bound the second term on the right-hand side of (4.5). Since $\Pi_{h}^{\nabla}v-E_{h}v\in V_{k}^{2,c}(K)$, by the inverse inequality and the norm equivalence in [21], one gets
+$|\Pi_{h}^{\nabla}v-E_{h}v|_{2,K}\lesssim h_{K}^{-2}\|\Pi_{h}^{\nabla}v-E_{h}v\|_{0,K}\lesssim h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\nabla}v-E_{h}v)\|_{l^{2}},\quad K\in\mathcal{T}_{h}.$
 where $\bm{\chi}^{c}$ is the set of DoFs on $K$ for the $C^{1}$-continuous space.
 In the following, we only provide the details of bounding the values and its gradient values of the function $\Pi_{h}^{\Delta}v-E_{h}v$ at vertices since the argument implies the treatment for moments.
-(1) Let $z\in\mathcal{V}_{h}^{0}$ be an interior vertex.
-Assume that $K_{1}=K,K_{2},\cdots,K_{L}$ share the node $z$, and denote $(\Pi_{h}^{\Delta}\phi)_{i}=\Pi_{h}^{\Delta}\phi|_{K_{i}}$, where $K_{i}$ and $K_{i+1}$ are two neighboring elements.
+Let $z\in\mathcal{V}_{h}^{0}$ be an interior vertex.
+Assume that $K_{1}=K,K_{2},\cdots,K_{L}$ share the node $z$, and denote $(\Pi_{h}^{\nabla}\phi)_{i}=\Pi_{h}^{\nabla}\phi|_{K_{i}}$, where $K_{i}$ and $K_{i+1}$ are two neighboring elements.
 We have
-$(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
-$=(\Pi_{h}^{\Delta}v)_{1}(z)-\frac{1}{L}((\Pi_{h}^{\Delta}v)_{1}+\cdots+(\Pi_{h}^{\Delta}v)_{L})(z)$
-$=\frac{1}{L}(((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2})+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{3})+\cdots+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{L}))(z).$
+$(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)$
+$=(\Pi_{h}^{\nabla}v)_{1}(z)-\frac{1}{L}((\Pi_{h}^{\nabla}v)_{1}+\cdots+(\Pi_{h}^{\nabla}v)_{L})(z)$
+$=\frac{1}{L}(((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2})+((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{3})+\cdots+((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{L}))(z).$
 Since $L$ is uniformly bounded and
-$(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{j}(z)=\sum\limits_{i=1}^{j-1}((\Pi_{h}^{\Delta}v)_{i}-(\Pi_{h}^{\Delta}v)_{i+1})(z),$
-it suffices to consider the term $(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)$. Apply the inverse inequality of polynomials to get
-$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|\leq\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{\infty,e}\lesssim h_{e}^{-1/2}\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}.$ (4.3)
+$(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{j}(z)=\sum\limits_{i=1}^{j-1}((\Pi_{h}^{\nabla}v)_{i}-(\Pi_{h}^{\nabla}v)_{i+1})(z),$
+it suffices to consider the term $(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{2}(z)$. Apply the inverse inequality of polynomials to get
+$|(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{2}(z)|\leq\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{\infty,e}\lesssim h_{e}^{-1/2}\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}.$ (4.6)
 Noting that $v$ is not continuous across $e$, we denote by $v_{i}$ the restriction of $v$ on $K_{i}$. Then one has
-$\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$ $=\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}+v_{1}-v_{2}+v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$
-$\leq\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}\|_{0,e}+\|v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}+\|[v]\|_{0,e}.$
+$\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}$ $=\|(\Pi_{h}^{\nabla}v)_{1}-v_{1}+v_{1}-v_{2}+v_{2}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}$
+$\leq\|(\Pi_{h}^{\nabla}v)_{1}-v_{1}\|_{0,e}+\|v_{2}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}+\|[v]\|_{0,e}.$
 By the trace inequality (2.5),
-$\|v-\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1/2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
+$\|v-\Pi_{h}^{\nabla}v\|_{0,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1/2}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
 If $z$ is a vertex on the domain boundary, then $E_{h}(v)(z)=0=v(z)$, which gives
-$|(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|=|\Pi_{h}^{\Delta}v(z)-v(z)|\leq\|\Pi_{h}^{\Delta}v-v\|_{\infty,K}.$
-By the Sobolev inequality (2.6),
-$\|v-\Pi_{h}^{\Delta}v\|_{\infty,K}\leq h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
-Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
-$\Big{(}h_{K}^{-1}(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2},$
+$|(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)|=|\Pi_{h}^{\nabla}v(z)-v(z)|\leq\|\Pi_{h}^{\nabla}v-v\|_{\infty,K}.$
+Using the Sobolev inequality (2.6) and the inverse inequality (3.6), we get
+$\|v-\Pi_{h}^{\nabla}v\|_{\infty,K}$ $\leq h_{K}|v-\Pi_{h}^{\nabla}v|_{2,K}+|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}$
+$\lesssim|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
+Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using norm equivalence (3.20) and (3.21), we obtain
+$\Big{(}h_{K}^{-1}(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e_{i}}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2},$
 where $e_{i}$ is the edge sharing by $K_{i}$ and $K_{i+1}$.
-(2) For the gradient value at an interior vertex $z\in\mathcal{V}^{0}$, one has
-$h_{z}\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
-$=$ $h_{z}(\nabla\Pi_{h}^{\Delta}v)_{1}(z)-\frac{h_{z}}{L}((\nabla\Pi_{h}^{\Delta}v)_{1}+\cdots+(\nabla\Pi_{h}^{\Delta}v)_{L})(z)$
-$=$ $\frac{h_{z}}{L}(((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{2})+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{3})+\cdots+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{L}))(z)$
-and we only need to consider the term $h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$. By the inverse inequality of polynomials,
-$h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$ $\lesssim h_{z}h_{e}^{-1/2}|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}$
-$\lesssim h_{e}^{1/2}\Big{(}|(\Pi_{h}^{\Delta}v)_{1}-v_{1}|_{1,e}+|v_{2}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}+|[v]|_{1,e}\Big{)}.$
-Using the trace inequality, we get
-$|v-\Pi_{h}^{\Delta}v|_{1,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{2,K}+h_{K}^{-1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}.$
+For the gradient value at an interior vertex $z\in\mathcal{V}^{0}$, one has
+$h_{z}\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)$
+$=$ $h_{z}(\nabla\Pi_{h}^{\nabla}v)_{1}(z)-\frac{h_{z}}{L}((\nabla\Pi_{h}^{\nabla}v)_{1}+\cdots+(\nabla\Pi_{h}^{\nabla}v)_{L})(z)$
+$=$ $\frac{h_{z}}{L}(((\nabla\Pi_{h}^{\nabla}v)_{1}-(\nabla\Pi_{h}^{\nabla}v)_{2})+\cdots+((\nabla\Pi_{h}^{\nabla}v)_{1}-(\nabla\Pi_{h}^{\nabla}v)_{L}))(z)$
+and we only need to consider the term $h_{z}((\nabla\Pi_{h}^{\nabla}v)_{1}(z)-(\nabla\Pi_{h}^{\nabla}v)_{2}(z))$. By the inverse inequality of polynomials,
+$h_{z}((\nabla\Pi_{h}^{\nabla}v)_{1}(z)-(\nabla\Pi_{h}^{\nabla}v)_{2}(z))\lesssim$ $h_{z}h_{e}^{-1/2}|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}|_{1,e}$
+$\lesssim$ $h_{e}^{1/2}(|(\Pi_{h}^{\nabla}v)_{1}-v_{1}|_{1,e}+|v_{2}-(\Pi_{h}^{\nabla}v)_{2}|_{1,e}+|[v]|_{1,e}).$
+Using the trace inequality (2.5) and inverse inequality (3.6), we get
+$|v-\Pi_{h}^{\nabla}v|_{1,e}\lesssim h_{K}^{-1/2}|v-\Pi_{h}^{\nabla}v|_{1,K}.$
 For the jump term $|[v]|_{1,e}$, noting that $[\partial_{\bm{t}_{e}}v]$ is a polynomial on $e$, we obtain from the inverse inequality for polynomials to get $\|[\partial_{\bm{t}_{e}}v]\|_{e}\lesssim h_{e}^{-1}\|[v]\|_{e}$ and hence
-$|[v]|_{1,e}\lesssim\|[\partial_{\bm{n}_{e}}v]\|_{0,e}+h_{e}^{-1}\|[v]\|_{e}.$
+$|[v]|_{1,e}\lesssim\|[\partial_{\bm{n}_{e}}v]\|_{0,e}+h_{e}^{-1}\|[v]\|_{0,e}.$
 For a boundary vertex $z$, we have $\nabla E_{h}(v)(z)=\nabla v(z)=0$. Let $e$ be the associated boundary edge. By the inverse inequality for polynomials and the trace inequality,
-$h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|$ $=h_{z}|\nabla\Pi_{h}^{\Delta}v(z)|\leq h_{K}\|\nabla\Pi_{h}^{\Delta}v\|_{\infty,e}\lesssim h_{e}^{1/2}\|\nabla\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v\|_{0,e}$
-$=h_{e}^{-1/2}\|\Pi_{h}^{\Delta}v-v\|_{0,e}\lesssim|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
-Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
-$\Big{(}h_{K}^{-1}h_{z}|\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\Big{(}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2}+\frac{1}{|e_{i}|}\|[\partial_{\bm{n}_{e_{i}}}v]\|_{0,e_{i}}^{2}\Big{)}.$
-This completes the proof. ∎
+$h_{z}|\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)|$ $=h_{z}|\nabla\Pi_{h}^{\nabla}v(z)|\leq h_{K}\|\nabla\Pi_{h}^{\nabla}v\|_{\infty,e}\lesssim h_{e}^{-1/2}\|\Pi_{h}^{\nabla}v\|_{0,e}$
+$=h_{e}^{-1/2}\|\Pi_{h}^{\nabla}v-v\|_{0,e}\lesssim|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
+Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using norm equivalence (3.20) and (3.21), we have
+$\Big{(}h_{K}^{-1}h_{z}|\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim\sum_{e_{i}}\sum_{K\in\mathcal{T}_{e_{i}}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e_{i}}\frac{1}{|e_{i}|}\|[\partial_{\bm{n}_{e_{i}}}v]\|_{0,e_{i}}^{2}.$
+The above arguments yield the first estimate (4.2).
+Step 2: Let $e$ be an edge shared by two elements $K^{-}$ and $K^{+}$. By the trace inequality (2.5) and the inverse inequality (3.6), one gets
+$\|[v]\|_{0,e}=\|[v-I_{h}^{c}v]\|_{0,e}\lesssim\sum_{K=K^{-},K^{+}}h_{K}^{-1/2}\|v-I_{h}^{c}v\|_{0,K}.$
+Noting that $\bm{\chi}(v)=\bm{\chi}(I_{h}^{c}v)$, we apply the norm equivalence (3.21) and the norm equivalence for $H^{1}$-conforming virtual elements (cf. [21]) to get
+$\|v-I_{h}^{c}v\|_{0,K}\leq\|v-\Pi_{h}^{\nabla}v\|_{0,K}+\|\Pi_{h}^{\nabla}v-I_{h}^{c}v\|_{0,K}\lesssim h_{K}\|\bm{\chi}(v-\Pi_{h}^{\nabla}v)\|_{\ell^{2}},$
+which yields the second estimate (4.3).
+Step 3: The third estimate follows by considering the triangle inequality
+$\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}\leq\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}+\Big{\|}\Big{[}\frac{\partial(v-\Pi_{h}^{\nabla}v)}{\partial\bm{n}_{e}}\Big{]}\Big{\|},$ (4.7)
+with the details omitted. ∎
 Remark 4.1.
-Our result aligns with the finite element methods since the first term on the right-hand side of (4.2) vanishes in this case (see [9, 8, 12] for instance).
-For the medius error analysis, we should bound the (semi-)norm of $v-E_{h}v$ by $|v|_{2,h}$.
-Here, we present the result for $\|v-E_{h}v\|_{0,K}$ as an example, even though this estimate will not be needed in the subsequent analysis.
-Lemma 4.2.
-For any $v\in V_{h}$ and $E_{h}v$ defined above, there holds
-$\sum\limits_{K\in\mathcal{T}_{h}}h_{K}^{-4}\|v-E_{h}v\|_{0,K}^{2}\lesssim|v|_{2,h}^{2}.$
-Proof.
-From the triangle inequality
-$\|v-E_{h}v\|_{0,K}\leq\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$
-and the standard estimate
-$h_{K}^{-2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\lesssim|v|_{2,K},$
-it remains to bound the term $\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}$. By the norm equivalence,
-$h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\simeq h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}}.$
-Let $z$ be an interior vertex.
-As in the proof of Lemma 4.1, one only needs to consider $|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$.
-Noting that $v$ is continuous at $z$ (i.e., $v_{1}(z)=v_{2}(z)$), by applying the Sobolev inequality (2.6), we obtain
-$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$ $\leq|(\Pi_{h}^{\Delta}v)_{1}(z)-v_{1}(z)|+|v_{2}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|$
-$\lesssim\sum_{K=K_{1},K_{2}}\Big{(}h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}\Big{)}.$
-The proof is finished by using the error estimate for $\Pi_{h}^{\Delta}$. ∎
+Our result aligns with the traditional finite element methods on a triangular mesh since the first term on the right-hand side of (4.2) vanishes when $v$ is replaced by a polynomial of degree $\leq k$ and the remaining parts are the jumps accounting for the discontinuities (see [9, 8, 12] for instance).
+We remark that the second estimate (4.3) implies that the interior penalty virtual element can be treated as a continuous element in the a posteriori estimate.
+This property is referred to as the quasi-$C^{0}$ continuity in this article.
 5 A posteriori error estimation
@@
 By reviewing the calculation in Section 5.1 of [16], we can incorporate the computable components of the local error estimator $\eta_{K}$ as follows:
-$\eta_{0,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{0,e}^{2}\Big{)}^{1/2},\qquad\eta_{0,e}=\frac{\lambda_{e}^{3/2}}{|e|^{3/2}}\|[u_{h}]\|_{0,e},$
-$\eta_{1,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2},\qquad\eta_{1,e}=\frac{\lambda_{e}^{1/2}}{|e|^{1/2}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e},$
-$\eta_{2,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{2,e}^{2}\Big{)}^{1/2},\qquad\eta_{2,e}=|e|^{1/2}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
-$\eta_{3,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2},\qquad\eta_{3,e}=|e|^{3/2}\Big{\|}\Big{[}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}}+\frac{\partial^{3}(\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
-$\eta_{4,K}=\big{(}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})\big{)}^{1/2},$
+$\eta_{1,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2},\qquad\eta_{1,e}=\frac{1}{|e|^{1/2}}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e},$
+$\eta_{2,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{2,e}^{2}\Big{)}^{1/2},\qquad\eta_{2,e}=|e|^{1/2}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
+$\eta_{3,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2},\qquad\eta_{3,e}=|e|^{3/2}\Big{\|}\Big{[}\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}_{e}}+\frac{\partial^{3}(\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e},$
+$\eta_{4,K}=h_{K}^{-1}\big{(}S^{K}(u_{h}-\Pi_{h}^{\nabla}u_{h},u_{h}-\Pi_{h}^{\nabla}u_{h})\big{)}^{1/2},$
 $\eta_{5,K}:=h_{K}^{2}\|f-f_{h}\|_{0,K},$
-$\eta_{6,K}=h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K},$ (5.1)
+$\eta_{6,K}=h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h}\|_{0,K},$ (5.1)
 where $f_{h}|_{K}=\Pi_{0,K}^{k}f$ and $\omega(K)$ denotes the union of all elements in $\mathcal{T}_{h}$ sharing some edges or points with $K$.
@@
 The local and global error estimators are respectively defined by
-$\eta_{K}(u_{h})=\Big{(}\sum\limits_{i=0}^{6}\eta_{i,K}^{2}\Big{)}^{1/2}$
-and
-$\eta(u_{h})=\eta(u_{h},\mathcal{T}_{h})=:\Big{(}\sum\limits_{K\in\mathcal{T}_{h}}\eta_{K}^{2}(u_{h})\Big{)}^{1/2}.$
+$\eta_{K}(u_{h})=\Big{(}\sum\limits_{i=1}^{6}\eta_{i,K}^{2}\Big{)}^{1/2},\quad\eta(u_{h})=\eta(u_{h},\mathcal{T}_{h})=\Big{(}\sum\limits_{K\in\mathcal{T}_{h}}\eta_{K}^{2}(u_{h})\Big{)}^{1/2}.$
+Note that we do not include $\eta_{0,K}$ in the error estimator, where $\eta_{0,K}$ is defined in (3.15).
 5.1   Upper bound
@@
 $\|e_{h}\|_{h}$ $\leq|u-E_{h}u_{h}|_{2}+|E_{h}u_{h}-u_{h}|_{2,h}+J_{1}(e_{h},e_{h})^{1/2}$ (5.2)
-$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},u_{h}-\Pi_{h}^{\Delta}u_{h})$
-$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|^{3}}\|[u_{h}]\|_{0,e}^{2}+\sum\limits_{e\in\mathcal{E}_{h}}\frac{1}{h_{e}}\Big{\|}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}+J_{1}(e_{h},e_{h})^{1/2}$
-$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}.$
+$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\nabla}u_{h},u_{h}-\Pi_{h}^{\nabla}u_{h})$
+$\qquad+\sum\limits_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}+J_{1}(e_{h},e_{h})^{1/2}$
+$\lesssim|u-E_{h}u_{h}|_{2}+\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}.$
 Therefore, we only need to consider the first term $|u-E_{h}u_{h}|_{2}$ in the last equation.
-Let $\phi=u-E_{h}u_{h}$ and denote by $I_{h}:H^{2}(\Omega)\to V_{h}$ the interpolation operator. According to the definitions of the continuous and discrete variational problems, one gets
+Let $\phi=u-E_{h}u_{h}$. According to the definitions of the continuous and discrete variational problems, one gets
 $|u-E_{h}u_{h}|_{2}^{2}=a(u-E_{h}u_{h},\phi)$
@@
 $I_{4}=(f-f_{h},\phi).$
-For the term $I_{1}$, using the Cauchy-Schwarz inequality and the estimate (4.2), one gets
+For the term $I_{1}$, using the Cauchy-Schwarz inequality and the estimate in Lemma 4.1, one gets
 $I_{1}$ $\leq\sum_{K\in\mathcal{T}_{h}}|u_{h}-E_{h}u_{h}|_{2,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-E_{h}u_{h}|_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}$
-$\lesssim\Big{(}\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{0,K}^{2}+\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
-For the term $I_{2}$, using the integration by parts we can split it into four terms as
-$I_{2}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}$
-$\qquad+\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}},\phi-I_{h}\phi\Big{)}_{\partial K}$
-$\qquad-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
-$\qquad+\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)$
-$=:I_{2,1}+I_{2,2}+I_{2,3}+I_{2,4}.$
-For $I_{2,1}$, by the interpolation error estimate,
-$I_{2,1}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)_{K}\leq\sum_{K\in\mathcal{T}_{h}}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}\|\phi-I_{h}\phi\|_{0,K}$
-$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h}\|_{0,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{6,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
-For the sake of brevity, we set $Q_{3}(u_{h})=\frac{\partial(\Delta\Pi_{h}^{\Delta}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}}$ in $I_{2,2}$.
-From the equality (2.2), one gets
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{1,K}^{2}+\eta_{4,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
+For the term $I_{2}$, using the integration by parts we can split it into several terms as
+$I_{2}$ $=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)+(f_{h},\phi-I_{h}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)$
+$=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)$
+$\qquad+\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)_{K}$
+$\qquad+\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}},\phi-I_{h}\phi\Big{)}_{\partial K}$
+$\qquad-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
+$\qquad+\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)$
+$=:I_{2,0}+I_{2,1}+I_{2,2}+I_{2,3}+I_{2,4}.$
+For $I_{2,0}$, we obtain from the boundedness of $\Pi_{h}^{\Delta}$ in Lemma 3.3, the interpolation error estimate (3.8), the inverse inequality (3.6) and the norm equivalence (3.20) that
+$I_{2,0}$ $=\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h},\phi-I_{h}\phi)\leq\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|\phi-I_{h}\phi|_{2,K}$
+$\lesssim\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}|\phi|_{2,K}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}|\phi|_{2,K}$
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
+For $I_{2,1}$, by the interpolation error estimate (3.8),
+$I_{2,1}$ $=\sum_{K\in\mathcal{T}_{h}}(f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h},\phi-I_{h}\phi)_{K}\leq\sum_{K\in\mathcal{T}_{h}}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h}\|_{0,K}\|\phi-I_{h}\phi\|_{0,K}$
+$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}u_{h}\|_{0,K}|\phi|_{2,K}\leq\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{6,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
+For the sake of brevity, we set $Q_{3}(u_{h})=\frac{\partial(\Delta\Pi_{h}^{\nabla}u_{h})}{\partial\bm{n}}+\frac{\partial^{3}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}\partial\bm{t}^{2}}$ in $I_{2,2}$.
+From Eq.
+(2.2) one gets
 $I_{2,2}=$ $\sum_{K\in\mathcal{T}_{h}}(Q_{3}(u_{h}),\phi-I_{h}\phi)_{\partial K}$
@@
 which indicates
-$I_{2,4}=\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)=0.$
-Hence, it follows from the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates that
+$I_{2,4}=\sum_{K\in\mathcal{T}_{h}}\sum\limits_{z\in\mathcal{V}(K)}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{t}\partial\bm{n}}\Big{]}(z)(\phi-I_{h}\phi)(z)=0.$
+Hence, it follows from the Cauchy-Schwarz inequality, the trace inequality (2.5) and the interpolation error estimates (3.8) that
 $I_{2,2}$ $=\sum\limits_{e\in\mathcal{E}_{h}^{0}}([Q_{3}(u_{h})],\{\phi-I_{h}\phi\})_{e}\leq\sum\limits_{e\in\mathcal{E}_{h}^{0}}\|[Q_{3}(u_{h})]\|_{0,e}\|\{\phi-I_{h}\phi\}\|_{0,e}$
@@
 $\lesssim\sum\limits_{e\in\mathcal{E}_{h}^{0}}h_{e}^{3/2}\|[Q_{3}(u_{h})]\|_{0,e}|\phi|_{2,\omega(e)}\lesssim\Big{(}\sum_{e\in\mathcal{E}_{h}^{0}}\eta_{3,e}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{3,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
-For $I_{2,3}$, noting that $[\frac{\partial\phi}{\partial\bm{n}}]_{e}=0$, from (2.2), we have
-$I_{2,3}$ $=-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
-$=-\sum_{e\in\mathcal{E}_{h}^{0}}\Big{(}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-\sum_{e\in\mathcal{E}_{h}}(\{\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}\},[\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}])_{e}\Big{)}$
-$=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\Delta}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-J_{2}(u_{h},I_{h}\phi).$
-Similarly, by using the Cauchy-Schwarz inequality, the trace inequality and the interpolation error estimates, we can get
-$I_{2,3}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
+For $I_{2,3}$, noting that $[\frac{\partial\phi}{\partial\bm{n}}]_{e}=0$, from (2.2) and the definition of $V_{k}(K)$, we have
+$I_{2,3}$ $=-\sum_{K\in\mathcal{T}_{h}}\Big{(}\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}},\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\Big{)}_{\partial K}$
+$=-\sum_{e\in\mathcal{E}_{h}^{0}}\Big{(}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-\sum_{e\in\mathcal{E}_{h}}(\{\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}\},[\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}])_{e}\Big{)}$
+$=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}-J_{2}(u_{h},I_{h}\phi),$ (5.5)
+which gives
+$I_{2,3}+J_{2}(u_{h},I_{h}\phi)=-\sum_{e\in\mathcal{E}_{h}^{0}}([\frac{\partial^{2}\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}^{2}}],\{\frac{\partial(\phi-I_{h}\phi)}{\partial\bm{n}}\})_{e}.$
+Similarly, by using the Cauchy-Schwarz inequality, the trace inequality (2.5) and the interpolation error estimates (3.8), we can get
+$I_{2,3}+J_{2}(u_{h},I_{h}\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
 Collecting the above estimates to derive
-$I_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{2,K}^{2}+\eta_{3,K}^{2}+\eta_{6,K}^{2})\Big{)}^{1/2}|\phi|_{2}-J_{2}(u_{h},I_{h}\phi).$
-For the term $I_{3}$, by the norm equivalence, we obtain
+$I_{2}+J_{2}(u_{h},I_{h}\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}(\eta_{2,K}^{2}+\eta_{3,K}^{2}+\eta_{4,K}^{2}+\eta_{6,K}^{2})\Big{)}^{1/2}|\phi|_{2}.$
+For the term $I_{3}$, by the norm equivalence (3.19) and (3.20), the boundedness of $\Pi_{h}^{\Delta}$ in Lemma 3.3, the inverse inequality (3.6), we obtain
 $I_{3}$ $=-\sum_{K\in\mathcal{T}_{h}}a^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\phi)+\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},I_{h}\phi-\Pi_{h}^{\Delta}I_{h}\phi)$
@@
 $\lesssim\sum_{K\in\mathcal{T}_{h}}\Big{(}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|{I_{h}}\phi-\Pi_{h}^{\Delta}({I_{h}}\phi)|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}|\phi|_{2,K}\Big{)}$
-$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}\Big{)}^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}\big{(}|\phi|_{2,K}^{2}+|\phi-{I_{h}}\phi|_{2,K}^{2}\big{)}\Big{)}^{1/2}$
-$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}\Big{)}^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}\big{(}|\phi|_{2,K}^{2}+|\phi-{I_{h}}\phi|_{2,K}^{2}\big{)}\Big{)}^{1/2}.$
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}+|\Pi_{h}^{\nabla}u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}\Big{)}^{1/2}|\phi|_{2}.$
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
 For $I_{4}$, it is easy to get
 $I_{4}=(f-f_{h},\phi)\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{5,K}^{2}(u_{h})\Big{)}^{1/2}|\phi|_{2}.$
-For $J_{1}(u_{h},I_{h}\phi)$, by the interpolation error estimate, we have
-$J_{1}(u_{h},I_{h}\phi)\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{e\in\mathcal{E}_{h}}|e|^{-1}\Big{\|}\Big{[}\frac{\partial(I_{h}\phi-\phi)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
-For $J_{3}(u_{h},I_{h}\phi)$, the boundedness of $\Pi_{h}^{\Delta}$ gives
-$J_{3}(u_{h},I_{h}\phi)$ $=\Big{|}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}I_{h}\phi}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{|}\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{2,e}^{2})^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\Delta}\phi|_{2,K}^{2}\Big{)}^{1/2}$
-$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{2,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$ (5.5)
+For $J_{1}(u_{h},I_{h}\phi)$, by the trace inequality (2.5), the boundedness of $\Pi_{h}^{\nabla}$ and interpolation error estimate (3.8), we have
+$J_{1}(u_{h},I_{h}\phi)$ $\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{e\in\mathcal{E}_{h}}|e|^{-1}\Big{\|}\Big{[}\frac{\partial(\Pi_{h}^{\nabla}I_{h}\phi-\phi)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}^{1/2}\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$
+For $J_{3}(u_{h},I_{h}\phi)$, by the Cauchy-Schwarz inequality, trace inequality (2.5), the inverse inequalities of polynomials and the boundedness of $\Pi_{h}^{\nabla}$, ones give
+$J_{3}(u_{h},I_{h}\phi)$ $=\Big{|}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}I_{h}\phi}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{|}\lesssim(\sum_{e\in\mathcal{E}_{h}}\eta_{1,e}^{2})^{1/2}\Big{(}\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}I_{h}\phi|_{2,K}^{2}\Big{)}^{1/2}$
+$\lesssim\Big{(}\sum_{K\in\mathcal{T}_{h}}\eta_{1,K}^{2}\Big{)}^{1/2}|\phi|_{2}.$ (5.6)
 The combination of the above estimates yields
@@
 By means of the technique of bubble functions frequently used in the a posteriori error analysis [21, 24], we can also derive the following lower bound of the a posteriori error estimator.
+The proof is similar to [21], but for the sake of completeness, we still provide a detailed proof.
 Theorem 5.2.
 Let $u\in V$ be the solution of (2.3). Then, for any $w_{h}\in V_{h}$, $K\in\mathcal{T}_{h}$ and $e\in\mathcal{E}(K)\cap\mathcal{E}_{h}^{0}$, there hold
-$h^{4}_{K}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}w_{h}\|_{0,K}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,K}^{2}+\eta_{5,K}^{2},$ (5.6)
-$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2},$ (5.7)
-$h_{e}^{3}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}}+\frac{\partial^{3}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2}.$ (5.8)
+$h^{4}_{K}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}w_{h}\|_{0,K}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,K}^{2}+\eta_{5,K}^{2},$ (5.7)
+$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2},$ (5.8)
+$h_{e}^{3}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}}+\frac{\partial^{3}\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2}.$ (5.9)
 Proof.
-We divide the proof into three steps. For brevity, we omit the subscript $e$ in $\bm{t}_{e}$ and $\bm{n}_{e}$. For the sake of brevity, we denote $v=\Pi_{h}^{\Delta}w_{h}$ in what follows.
+We divide the proof into three steps. For brevity, we omit the subscript $e$ in $\bm{t}_{e}$ and $\bm{n}_{e}$. For the sake of brevity, we denote $v=\Pi_{h}^{\nabla}w_{h}$ in what follows.
 Step 1: For any $K\in\mathcal{T}_{h}$, assume $\mathcal{T}_{K}$ is the auxiliary triangulation of $K$, and let $b_{K}=\sum_{\tau\in\mathcal{T}_{K}}(\lambda_{1,\tau}\lambda_{2,\tau}\lambda_{3,\tau})^{2}$ be the element bubble function defined on $K$, where $\lambda_{i,\tau}$($i=1,2,3$) are the nodal basis functions at vertices of virtual triangles $\tau\in\mathcal{T}_{K}$.
@@
 According to the technique of bubble functions, we need to construct an auxiliary polynomial $\phi\in H^{2}_{0}(T)$ satisfying $\partial_{\bm{n}}\phi|_{\partial{T}\setminus{e}}=0$ and
-$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{e}$ (5.9)
+$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{e}.$ (5.10)
 Since the jump $[\frac{\partial^{2}v}{\partial\bm{n}^{2}}]|_{e}$ is a polynomial on $e$, we extend it to $T$, denoted by $E_{T}([\frac{\partial^{2}v}{\partial\bm{n}^{2}}])$, so that this function is constant along the lines perpendicular to $e$.
@@
 Since $\phi\in H_{0}^{2}(T)$, the equation (5.12) holds. Then, using (5.12), (5.15), the Cauchy-Schwarz inequality and the inverse inequalities of polynomials, we have
-$\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim\sum\limits_{\tau=\tau^{+},\tau^{-}}\Big{(}(\Delta^{2}v-f,\,\phi)_{\tau}+a^{\tau}(u-v,\,\phi)\Big{)}+\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{0,e}$
+[…]
 $\lesssim\sum_{\tau=\tau^{+},\tau^{-}}\Big{(}\|f-\Delta^{2}v\|_{0,\tau}+h_{\tau}^{-2}|v-u|_{2,\tau}\Big{)}\|\phi\|_{0,\tau}+\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}_{0,e}\|\partial_{\bm{n}}\phi\|_{0,e}.$ (5.16)
@@
 It follows from the definition of $\eta(u_{h})$ and Theorem 5.2 that
-$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})$
+$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})$
 $+\sum_{K\in\mathcal{T}_{h}}\eta_{4,K}^{2}+\sum_{K\in\mathcal{T}_{h}}\eta_{5,K}^{2}.$ (5.20)
@@
 $\eta_{5,K}=h_{K}^{2}\|f-f_{h}\|_{0,K}\lesssim h^{k+2}\|f\|_{k-2}=\mathcal{O}(h^{k+2}).$ (5.21)
-By the norm equivalence (3.17),
-$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$ $\lesssim|u-u_{h}|_{2,K}^{2}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$
-$\lesssim|u-u_{h}|_{2,K}^{2}+S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})$
+By the inverse inequality (3.6) and the norm equivalence (3.20),
+$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}$ $\lesssim|u-u_{h}|_{2,K}^{2}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}$
+$\lesssim|u-u_{h}|_{2,K}^{2}+h_{K}^{-2}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}^{2}$
 $\lesssim|u-u_{h}|_{2,K}^{2}+\eta_{4,K}^{2}.$ (5.22)
-It follows from the norm equivalence in (3.17) and the triangle inequality that
-$\eta_{4,K}$ $=h_{K}^{-1}\|\bm{\chi}(u_{h}-\Pi_{h}^{\Delta}u_{h})\|_{\ell^{2}}\eqsim|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}\lesssim|u-u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}.$
-Here, by the triangle inequality, (3.8) and (2.4), we derive
-$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|\Pi_{h}^{\Delta}u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u|_{2,K}\lesssim|u-u_{h}|_{2,K}+h^{k-1}.$
-Inserting (5.20) and (5.2) into (5.2), one has
-$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})+\mathcal{O}(h^{2k-2}).$ (5.22)
+It follows from the norm equivalence in (3.20), the inverse inequality (3.6) and the triangle inequality that
+$\eta_{4,K}$ $=h_{K}^{-1}\|\bm{\chi}(u_{h}-\Pi_{h}^{\nabla}u_{h})\|_{\ell^{2}}\eqsim h_{K}^{-1}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}\lesssim|u-u_{h}|_{2,K}+|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}.$
+Here, by the triangle inequality, Lemma 3.3 and (2.4), we derive
+$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}\leq|\Pi_{h}^{\nabla}u-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|u-\Pi_{h}^{\nabla}u|_{2,K}\lesssim|u-u_{h}|_{2,K}+h^{k-1}.$
+From the continuity of $u$, the trace inequality and (2.4), we get
+$J_{1}(u-u_{h},u-u_{h})$ $\lesssim\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial(\Pi_{h}^{\nabla}u-u)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}$
+$\lesssim\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}u-u|_{2,K}^{2}+h_{K}^{-2}|\Pi_{h}^{\nabla}u-u|_{1,K}^{2}+J_{1}(u_{h},u_{h})$
+$\lesssim h^{2k-2}+J_{1}(u_{h},u_{h}).$ (5.23)
+Inserting (5.21),(5.2) and (5.2) into (5.2), one has
+$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-u_{h}|_{2,K}^{2}+J_{1}(u_{h},u_{h})+\mathcal{O}(h^{2k-2}).$ (5.24)
 The proof is completed by the definition of the norm $\|\cdot\|_{h}$. ∎
+Remark 5.1.
+In the lower bound estimate, the error estimates for $u-\Pi_{h}^{\nabla}u$ are not the a priori estimates for the discrete problem.
 5.3   Comparison with the analysis of Morley-type virtual elements
@@
 See the proof of [21, Corollary 4.1]. ∎
-Our analysis of the upper bound in Theorem 5.1 can be extended to the Morley-type VEM. Utilizing Lemma 5.1, we can avoid the tedious integration by parts for the term $I_{2}$ as follows:
+Our analysis of the upper bound in Theorem 5.1 can be extended to the Morley-type VEM. Using Lemma 5.1, we can avoid tedious integration by parts for the term $I_{2}$ as follows:
 $I_{2}=(f_{h},\phi-I_{h}^{2}\phi)-\sum_{K\in\mathcal{T}_{h}}a^{K}(\Pi_{h}^{\Delta}u_{h},\phi-I_{h}^{2}\phi)=(f_{h},\phi-I_{h}^{2}\phi)=(f_{h}-\Delta^{2}\Pi_{h}^{\Delta}u_{h},\phi-I_{h}^{2}\phi),$
@@
 However, this does not apply to other types of VEMs for the biharmonic equation even for the $C^{1}$-continuous VEMs discussed in [16].
-Another difference in the error estimators from [13] is the substitution of the local estimators $\eta_{0,e}(u_{h})$ and $\eta_{1,e}(u_{h})$ with $\eta_{0,e}(\Pi_{h}^{\Delta}u_{h})$ and $\eta_{1,e}(\Pi_{h}^{\Delta}u_{h})$.
 For the Morley-type VEM in [13], unlike the triangle inequality
@@
 used in (5.2), they utilize the following inequality
-$|u-u_{h}|_{2,K}\leq|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-u_{h}|_{2,K}.$ (5.24)
+$|u-u_{h}|_{2,K}\leq|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-u_{h}|_{2,K}.$ (5.26)
 The second term corresponds to the error estimator $\eta_{4,K}$.
-For the first term, since $\Pi_{h}^{\Delta}u_{h}|_{K}$ is a polynomial on $K$, they treat $\Pi_{h}^{\Delta}u_{h}$ as a finite element function on the virtual triangulation $\widehat{\mathcal{T}}$, defining $\widehat{E}_{h}:\mathbb{P}_{2}(\widehat{\mathcal{T}})\to H_{0}^{2}(\Omega)$ as the enriching operator for the FEM.
+For the first term, since $\Pi_{h}^{\nabla}u_{h}|_{K}$ is a polynomial on $K$, they treat $\Pi_{h}^{\nabla}u_{h}$ as a finite element function on the virtual triangulation $\widehat{\mathcal{T}}$, defining $\widehat{E}_{h}:\mathbb{P}_{2}(\widehat{\mathcal{T}})\to H_{0}^{2}(\Omega)$ as the enriching operator for the FEM.
 They then decompose the first term as
-$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|u-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}.$ (5.25)
-Let $\widehat{E}_{h}\Pi_{h}^{\Delta}:=\widetilde{E}_{h}$. The decomposition given by (5.24) and (5.25) can be written as
+$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}\leq|u-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,K}.$ (5.27)
+Let $\widetilde{E}_{h}=\widehat{E}_{h}\Pi_{h}^{\nabla}$. The decomposition given by (5.26) and (5.27) can be written as
 $|u-u_{h}|_{2,K}$ $\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}$
-$\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}.$
-The first term can be estimated as for the first term in (5.23), namely the first term in (5.2) for Theorem 5.1. By using the standard result for $\widehat{E}$, the third term is bounded by
-$|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,h}^{2}=|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
-For any $v\in V_{h}$, it is evident that $E_{h}\phi$ can be defined for $\phi=\Pi_{h}^{\Delta}v$.
-By the definition of $E_{h}$, we have $E_{h}v=E_{h}\Pi_{h}^{\Delta}v$, which along with Lemma 4.1 implies
-$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
+$\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}.$
+The first term can be estimated as for the first term in (5.25), namely the first term in (5.2) for Theorem 5.1. By using the standard result for $\widehat{E}_{h}$, the third term is bounded by
+$|\Pi_{h}^{\nabla}u_{h}-\widetilde{E}_{h}u_{h}|_{2,h}^{2}=|\Pi_{h}^{\nabla}u_{h}-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\nabla}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
+For any $v\in V_{h}$, it is evident that $E_{h}\phi$ can be defined for $\phi=\Pi_{h}^{\nabla}v$.
+By the definition of $E_{h}$, we have $E_{h}v=E_{h}\Pi_{h}^{\nabla}v$, which along with Lemma 4.1 implies
+$|\Pi_{h}^{\nabla}v-E_{h}v|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\nabla}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
 Therefore, we can perform the similar decomposition
 $|u-u_{h}|_{2,K}$ $\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-E_{h}u_{h}|_{2,K}$
-$\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-E_{h}u_{h}|_{2,K}.$
-This approach allows us to derive the error estimators in [13] for the Morley VEMs of arbitrary order, not just the lowest order.
-However, for interior penalty methods, this may not be appropriate, as including the elliptic projector $\Pi_{h}^{\Delta}$ in $J_{1}$ (see (3.11)) is not feasible because $\|\cdot\|_{h}$ may not be a norm on $V_{h}$ in this context and $J_{1}(u-u_{h},u-u_{h})$ will introduce the a-priori information.
+$\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-E_{h}u_{h}|_{2,K}.$
+This approach allows us to directly derive the error estimators in [13] for the Morley VEMs of arbitrary order.
 6 Numerical experiments
-In this section, we present several benchmark tests for which the exact solution is known or at least its behavior is known in a qualitative sense.
+In this section, we present several benchmark tests for which the exact solution is known.
 These tests illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM.
 Unless otherwise specified, we only consider the lowest-order element $k=2$.
-Since the VEM solution $u_{h}$ is not explicitly known inside the polygonal elements, we evaluate the projection error, which is quantified by
-[…]
-The first term in (3.9) will be disregarded when focusing solely on the $H^{2}$-semi-norm.
+We also provide the treatment of inhomogeneous boundary value conditions.
+6.1   Treatment of inhomogeneous boundary value conditions
+Consider the fourth-order problem with inhomogeneous boundary value conditions
+$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=g_{D},\quad\dfrac{\partial u}{\partial\bm{n}}=g_{N}&{\rm on}~{}~{}\partial\Omega,\end{array}\right.$ (6.1)
+where $g_{D}\in H^{3/2}(\partial\Omega)$ and $g_{N}\in H^{1/2}(\partial\Omega)$ are prescribed functions. Let $v\in H^{2}(K)$. The integration by parts gives
+$(f,v)_{K}=(\Delta^{2}u,v)_{K}=(\nabla^{2}u,\nabla^{2}v)_{K}$
+$\hskip 85.35826pt+\int_{\partial K}\Big{(}\frac{\partial\Delta u}{\partial\bm{n}_{K}}v+\frac{\partial^{2}u}{\partial\bm{n}_{K}\partial\bm{t}_{K}}\frac{\partial v}{\partial\bm{t}_{K}}\Big{)}\mathrm{d}s-\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v}{\partial\bm{n}_{K}}\mathrm{d}s.$
+If $v$ is $C^{0}$-continuous, the second term on the right-hand side naturally vanishes when we sum the equation over all polygons in $\mathcal{T}_{h}$.
+However, when $v=v_{h}$ is a function in the interior penalty virtual element space $V_{h}$, which is not continuous, the second term contributes to the summation.
+The IPVEM in [30] neglects this contribution as it introduces a consistency error with a desired convergence order as shown in [30, 18], which gives
+$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s.$
+Using the formula in (2.2) and noting that $u\in H^{3}(\Omega)$, we get
+$\sum_{K\in\mathcal{T}_{h}}\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v_{h}}{\partial\bm{n}_{K}}\mathrm{d}s$ $=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}^{0}}\int_{e}\Big{[}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\{}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{\}}\mathrm{d}s$
+$=\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
+yielding
+$(f,v_{h})\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
+To ensure the symmetry and enforce the $C^{1}$-continuity of the discrete variational problem, we rewrite it as
+$(f,v_{h})+\sum_{e\in\mathcal{E}_{h}^{\partial}}\Big{(}-\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{)}$
+$\approx(\nabla^{2}u,\nabla^{2}v_{h})-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}u}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s$
+$\qquad+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
+where we have used the fact that $[u]_{e}=0$ and $[\frac{\partial u}{\partial\bm{n}_{e}}]_{e}=0$ for $e\in\mathcal{E}_{h}^{0}$.
+$\lambda_{e}\geq 1$ is some edge-dependent parameter.
+Taking into account Remark 3.1, we can describe the IPVEM for (6.1) with inhomogeneous boundary value conditions as follows: Find $u_{h}\in V_{h}^{g}=\{v\in V_{h}:v=I_{h}g_{D}~{}{\rm on}~{}\partial\Omega\}$ such that
+$\mathcal{A}_{h}(u_{h},v_{h})=F_{h}(v_{h}),\quad v_{h}\in V_{h},$ (6.2)
+where $g_{D}$ has been extended to the whole domain. The right-hand side is
+$F_{h}(v_{h})=\langle f_{h},v_{h}\rangle+\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{(}-g_{N}\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}+g_{N}\frac{\lambda_{e}}{|e|}\frac{\partial\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}}\Big{)}\mathrm{d}s.$
 Standard adaptive algorithms based on the local mesh refinement can be written as loops of the form
@@
 Our code is available from GitHub (https://github.com/Terenceyuyue/mVEM) as part of the mVEM package which contains efficient and easy-following codes for various VEMs published in the literature.
-The subroutine IPVEM_Zhao.m is used to compute the numerical solutions.
-The test script main_IPVEM_Zhao_avem.m verifies the convergence rates.
-The error estimator is computed by the subroutine IPVEM_Zhao_indicator.m.
+The subroutine IPVEM.m is used to compute the numerical solutions.
+The test script main_IPVEM_avem.m verifies the convergence rates.
+The error estimator is computed by the subroutine IPVEM_indicator.m.
+Since the VEM solution $u_{h}$ is not explicitly known inside the polygonal elements, we evaluate the projection error, which is quantified by
+[…]
 6.2   Solution without large gradients
@@
 $u(x,y)=xy(1-x)(1-y){\mathrm{exp}}\left(-1000((x-0.5)^{2}+(y-0.117)^{2})\right).$
-Fig. 4: The initial and the final adapted meshes. (a) The initial mesh; (b) After 20 refinement steps; (c) After 30 refinement steps; (d) The zoomed mesh in (c)
+Fig. 4: The initial and the final adapted meshes. (a) The initial mesh; (b) After 10 refinement steps; (c) After 20 refinement steps; (d) The zoomed mesh in (c)
 Fig. 5: Convergence rates of the error $|u-\Pi_{h}^{\Delta}u_{h}|_{2,h}$ and the error estimator $\eta(u_{h})$
@@
 (b) Numerical
-Fig. 6: The exact and numerical solutions
+Fig. 6: The exact and numerical solutions for Example 6.2
 For this example, we choose the parameter $\theta=0.6$.
-The initial mesh and the final adapted meshes after 20 and 30 refinement steps are presented in Fig.
+The initial mesh and the final adapted meshes after 10 and 20 refinement steps are presented in Fig.
 4 (a-c), respectively.
@@
 We plot the adaptive convergence order in Fig.
-6, which confirms the theoretical prediction as in the last example.
+5, which confirms the theoretical prediction as in the last example.
 We also display the numerical and exact solutions in Fig.
-5, from which we see the adaptive strategy correctly refines the mesh in a neighborhood of the singular point and there is a good level of agreement between the $H^{2}$ error and error estimator.
-7 Conclusion
-We developed a residual-type a posteriori error estimator for an interior penalty virtual element method applied to a Kirchhoff plate bending problem.
-A key step involves constructing an enriching operator to establish the reliability and efficiency of the a posteriori error bound.
-We presented an error estimate for the enriching operator that aligns with the interior penalty finite element method, as detailed in Lemma 4.1.
-Additionally, we provided an example estimate in Lemma 4.2 with a proof idea applicable to the medius error analysis.
+6, from which we see the adaptive strategy correctly refines the mesh in a neighborhood of the singular point and there is a good level of agreement between the $H^{2}$ error and error estimator.
+6.4   Solution with low regularity
+Fig. 7: Convergence rates of the error $|u-\Pi_{h}^{\Delta}u_{h}|_{2,h}$ and the error estimator $\eta(u_{h})$
+Example 6.3.
+This example examine the performance of the IPVEM under the low regularity condition.
+The domain is takes as an L-shape square $\Omega=(0,1)^{2}\backslash([1/2,1]\times[0,1/2])$.
+We select the load term $f$ such that the analytical solution is
+$u(x,y)=((x-0.5)^{2}+(y-0.5)^{2})^{5/6}.$
+It is evident that the solution $u\in H^{8/3-\epsilon}$ for some $\epsilon>0$ exhibits singular behavior at the point $(1/2,1/2)$.
+For error evaluation we introduce a small parameter $\delta$ into the exact solution.
+Specifically, we modify $u$ as $u(x,y)=((x-0.5)^{2}+(y-0.5)^{2}+\delta)^{5/6}$.
+In this example, we choose the parameter $\theta=0.6$.
+We plot the adaptive convergence order in Fig.
+7, from which we observe that the IPVEM achieves the correct convergence order, even under the low regularity condition of the solution.
+Additionally, we present the numerical and exact solutions in Fig.
+8, which demonstrate that the adaptive strategy effectively refines the mesh in the vicinity of the singular point.
+Fig. 8: The exact and numerical solutions for Example 6.3
+6.5   Formulation based on $H^{2}$-elliptic projector
+In [13, 16], the $H^{1}$-elliptic projector $\Pi_{h}^{\nabla}$ is replaced by the $H^{2}$-elliptic projector $\Pi_{h}^{\Delta}$ in the error estimators.
+For the IPVEM, we can adopt a similar approach, except for the handling of gradient jumps.
+In this case, we define
+$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
+$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
+$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
+The substitution of the $H^{1}$ projection in the gradient jumps creates difficulty in the process of $I_{2,3}$ in (5.5), as we cannot cancel $J_{2}$.
+On the other hand, using the $H^{1}$ projection is the preferred choice in the a priori estimate in [18].
+We can still prove that $\|w\|_{h}=(|w|_{2,h}^{2}+J_{1}(w,w))^{1/2}$ introduces a norm on $V_{h}$.
+It is enough to prove that $\|v_{h}\|_{h}=0$ implies $v_{h}=0$ for any given $v_{h}\in V_{h}$.
+By definition, $\|v_{h}\|_{h}=0$ is equivalent to $|v_{h}|_{2,h}=0$ and $J_{1}(v_{h},v_{h})=0$.
+Equation $|v_{h}|_{2,h}=0$ shows that $\nabla_{h}v_{h}$ is a piecewise constant on $\mathcal{T}_{h}$.
+On the other hand, the direct manipulation yields
+$\int_{e}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s$ $\leq\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\Delta}v_{h}}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s+\int_{e}\Big{[}\frac{\partial(v_{h}-\Pi_{h}^{\Delta}v_{h})}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s=\int_{e}\Big{[}\frac{\partial(v_{h}-\Pi_{h}^{\Delta}v_{h})}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s,$
+$\leq C\sum_{K=K^{-},K^{+}}(|v_{h}-\Pi_{h}^{\Delta}v_{h}|_{1,K}+|v_{h}-\Pi_{h}^{\Delta}v_{h}|_{2,K})\leq C\sum_{K=K^{-},K^{+}}|v_{h}|_{2,K}=0.$
+Since $\nabla_{h}v_{h}$ is piecewise constant, we further obtain $[\frac{\partial v_{h}}{\partial\bm{n}_{e}}]=0$ over the edges of $\mathcal{T}_{h}$.
+That is, the normal derivative of $v_{h}$ is continuous at the interior edges and vanishes at the boundary of $\Omega$.
+This reduces to the discussion in the proof of Lemma 4.2 in [30], so we omit the remaining argument.
+We repeated the test for all the examples and obtained similar results, so we do not include them here.
+The test script main_IPVEM_avemH2.m verifies the convergence rates.
+The error estimator is computed by the subroutine IPVEM_indicatorH2.m.
 CRediT authorship contribution statement
-Fang Feng and Yue Yu collaborated closely to shape the conceptualization, methodology, and writing of this research. Additionally, Yue Yu took charge of implementing the discrete method in the study.
+All authors worked closely together on the conceptualization, methodology, and writing of this research.
+Additionally, Yue Yu and Yuming Hu were responsible for implementing the discrete method used in the study.
 Declaration of competing interest
@@
 Data availability
-Data sharing is not applicable to this article as no data was generated or analyzed.
+No data was used for the research described in the article.
 Acknowledgements
-Fang Feng was supported by the National Science Foundation for Young Scientists of China (No.
-12401528) and also partially supported by Central Government Special Fund for Basic Scientific Research Business Expenses of Colleges and Universities, No.
+Fang Feng was partially supported by the National Natural Science Foundation of China (NSFC) grant 12401528 and the Fundamental Research Funds for the Central Universities, No.
 30924010837.
 Yue Yu was partially supported by the National Science Foundation for Young Scientists of China (No.
-12301561).
+12301561) and the Key Project of Scientific Research Project of Hunan Provincial Department of Education (No.
+24A0100).
+Jikun Zhao was partially supported by National Natural Science Foundation of China (No.
+12371411).